### v1

# Limit Computation Over Posets via Minimal Initial Functors

Abstract.
It is well known that limits can be computed by restricting along an initial functor, and that this can often simplify limit computation. We systematically study the algorithmic implications of this idea for diagrams indexed by a finite poset. We say an initial functor $F\colon C\to D$ with $C$ small is minimal if the sets of objects and morphisms of $C$ each have minimum cardinality, among the sources of all initial functors with target $D$. For $Q$ a finite poset or $Q\subseteq\mathbb{N}^{d}$ an interval (i.e., a convex, connected subposet), we describe all minimal initial functors $F\colon P\to Q$ and in particular, show that $F$ is always a poset inclusion. We give efficient algorithms to compute a choice of $F$. In the case that $Q\subseteq\mathbb{N}^{d}$ is an interval, we give asymptotically optimal bounds on $|P|$, the number of relations in $P$ (including identities), in terms of the number $n$ of minima of $Q$: We show that $|P|=\Theta(n)$ for $d\leq 3$, and $|P|=\Theta(n^{2})$ for $d>3$. We apply these results to give new bounds on the cost of computing $\lim G$ for a functor $G\colon Q\to\mathbf{Vec}$ valued in vector spaces. For $Q$ connected, we also give new bounds on the cost of computing the generalized rank of $G$ (i.e., the rank of the induced map $\lim G\to\operatorname{colim}G$), which is of interest in topological data analysis.
1. Introduction
Let $\mathbf{Vec}$ denote the category of vector spaces over a fixed field $\mathbbm{k}$. Given a category $D$ with finitely many objects and morphisms and a functor $G\colon D\to\mathbf{Vec}$, a standard equalizer formula [mac2013categories, Theorem V.2.2] expresses $\lim G$ as a kernel of a matrix, which can be computed via Gaussian elimination. However, this matrix can be large, and reducing it can be expensive.
In general, one can compute the limit of a functor $G\colon D\to E$ by restricting along an initial functor $F\colon C\to D$; see Section˜2.3. If the category $C$ has far fewer objects and morphisms than $D$, then restricting along $F$ can substantially simplify the problem of computing $\lim G$. While this idea is well known, to our knowledge its algorithmic implications have not been systematically studied.
In this paper, we study these implications in the case that $D=Q$ is a poset and $E=\mathbf{Vec}$. (More generally, the core ideas of our approach apply when $E$ is any category where products and equalizers can be computed.) We develop theory and algorithms both for the case that $Q$ is a finite poset and the case that $Q$ is an interval in $\mathbb{N}^{d}$, (i.e., a connected, convex subposet; see Definition˜2.3). Here, $\mathbb{N}^{d}$ is given the product partial order, i.e., $(y_{1},\ldots,y_{d})\leq(z_{1},\ldots,z_{d})$ if and only if each $y_{i}\leq z_{i}$. Our motivation for considering intervals in $\mathbb{N}^{d}$ arises from computational questions in topological data analysis, and specifically in multiparameter persistent homology [carlsson2009theory, botnan2023introduction], where functors $\mathbb{N}^{d}\to\mathbf{Vec}$ are the central algebraic objects of study.
We develop a theory of minimal initial functors $F\colon P\to Q$, i.e., those for which the sets of objects and morphisms of $P$ both have minimum cardinality, among all sources of all initial functors. We give algorithms for computing such functors, and work out the consequences of these results for limit computation. Here, we summarize our main results, deferring the precise statements to Section˜3.
  1. (1)
    We present a structure theorem for minimal initial functors, Theorem˜3.5, which establishes the existence of a minimal initial functor $F\colon P\to Q$ and gives an explicit description of all such $F$. Specifically, the theorem says that, up to unique isomorphism, any such $F$ is the inclusion of a certain type of subposet $P\subseteq Q$, which we call an initial scaffold. While $Q$ can have multiple non-isomorphic initial scaffolds, all have the same underlying set of objects (but different sets of relations). As we discuss in Remark˜3.7, related results concerning full subposets have appeared in a recent paper of Brüstle, Desrochers, and Leblanc [brustle2025generalized].
  2. (2)
    For $Q\subseteq\mathbb{N}^{d}$ an interval, Theorem˜3.10 bounds $|P|$, the number of relations of $P$ (including identity relations), in terms of the number $n$ of minima of $Q$, as follows:
    $|P|=\begin{cases}\Theta(n)&\textup{ for }d\leq 3,\\ \Theta(n^{2})&\textup{ for }d>3.\end{cases}$
    To obtain these bounds, we establish a connection between initial scaffolds of intervals in $\mathbb{N}^{d}$ and the support of Betti numbers of monomial ideals. We then apply a well-known bound on the Betti numbers of monomial ideals, due to Bayer, Peeva, and Sturmfels [bayer1998monomial].
  3. (3)
    For $Q$ any finite poset, we give an algorithm to compute a choice of $F$ from the Hasse diagram of $Q$. We also give two specialized algorithms to compute $F$ when $Q\subseteq\mathbb{N}^{d}$ is an interval, one for the case $d\leq 3$ and one for arbitrary $d$. Theorems˜3.13 and 3.17 bound the complexity of each algorithm. Our algorithms for the interval case do not require the Hasse diagram of $Q$ as input, but instead take a more parsimonious representation of $Q$, which we call the upset presentation of $Q$; see Definition˜3.14. If the interval $Q\subseteq\mathbb{N}^{d}$ is finite, then we can instead specify $Q$ via its sets of minima and maxima.
  4. (4)
    We apply these results to derive new algorithms and bounds for computing the (co)limit of a functor $G\colon Q\to\mathbf{Vec}$ (Corollaries˜3.23 and 3.24). The statements of these bounds depend on how $G$ is input to the algorithms. For one concrete corollary of these bounds, suppose $Q\subseteq\mathbb{N}^{d}$ is a finite interval with a total of $s$ minima and maxima and we are given a free presentation $F_{1}\to F_{0}$ of $G$ where $F_{1}$ and $F_{0}$ have total rank $r$. Letting $\omega<2.373$ denote the exponent of matrix multiplication, Corollary˜3.31 implies that the cost of computing $\lim G$ is
    $O(s\log s+r^{3})$  $\quad\textup{ for }d=2,$
    $O((rs)^{\omega})$  $\quad\textup{ for }d=3,$
    $O(r^{\omega}s^{\omega+1})$  $\quad\textup{ for }d>3.$
    Our results on limit computation also encompass the case that $G$ is given implicitly as the homology of a chain complex of free functors, as well as the case that matrix representations of the structure maps of $G$ are explicitly given.
  5. (5)
    Given a connected, finite poset $Q$ and functor $G\colon Q\to\mathbf{Vec}$, we apply our results to bound the cost of computing the generalized rank of $G$, which is the rank of the induced map $\lim G\to\operatorname{colim}G$; see Corollaries˜3.36 and 3.37. As we discuss in Section˜3.5, this problem is of interest in TDA and has been studied in prior work [DKM22, dey2024computing, asashiba2024interval]. Our bounds extend and improve on prior bounds by Dey, Kim, and Mémoli [DKM22] and by Dey and Xin [dey2024computing].
Organization
In Section˜2, we cover preliminaries about posets, limits, initial functors, and $\mathbf{Vec}$-valued functors. Section˜3 is the heart of our narrative: Here, we precisely state the main definitions and results of our work, together with examples and context. The remaining sections of the paper give the algorithms and proofs underlying our main results: Section˜4 gives the proof of Theorem˜3.5, our structure theorem for minimal initial functors; Section˜5 gives the proof of our size bound Theorem˜3.10; Section˜6 presents our algorithms for computing initial scaffolds, and proves the bounds on their complexity; Section˜7 completes the proofs of our bounds on limit computation; and Section˜8 proves our bounds on generalized rank computation. Section˜9 briefly discussions directions for future work.
Acknowledgments
We thank Ezra Miller for pointing out that Lemma˜5.12 follows from a version of Hochster’s formula [miller2005combinatorial, Theorem 1.34]. TD acknowledges the support of NSF grants DMS 2301360, and CCF 2437030. ML acknowledges the support of a grant from the Simons Foundation (Award ID 963845).
2. Preliminaries
2.1. Categories and Posets
We assume familiarity with a few very basic notions in category theory, e.g., categories, (full) functors, and natural transformations. For an introduction to category theory, see, e.g., the textbook [riehl2017category, mac2013categories].
We denote the collections of objects and morphisms of a category $C$ as $\operatorname{Ob}C$ and $\hom C$, respectively. For $c,c^{\prime}\in\operatorname{Ob}C$, the set of morphisms from $c$ to $c^{\prime}$ is denoted $\hom(c,c^{\prime})$. We use subscripts to denote the action of a functor on objects and morphisms. That is, given a functor $F\colon C\to D$ and $c\in\operatorname{Ob}C$ we let $F_{c}=F(c)$, and given a morphism $g\colon c\to c^{\prime}$, we let $F_{g}=F(g)$. We call $C$ thin if $|\hom(c,c^{\prime})|\leq 1$ for all $c,c^{\prime}\in\operatorname{Ob}C$. If $C$ is thin, then $g\colon c\to c^{\prime}$ is determined by its source and target, so we write $F_{cc^{\prime}}=F(g)$.
Definition 2.1.
A poset is a pair $(Q,\leq)$, where $Q$ is a set and $\leq$ is a binary relation on $Q$, called the partial order, satisfying the following properties.
  • $q\leq q$ for all $q\in Q$,
  • if $p\leq q$ and $q\leq p$, then $p=q$,
  • if $p\leq q$ and $q\leq r$, then $p\leq r$.
When $p\leq q$ and $p\neq q$, we write $p<q$. We often abuse notation slightly and let $Q$ denote the poset $(Q,\leq)$. By a further slight abuse, the elements of $\leq$ are called relations.
We regard a poset $Q$ as a thin category where $\operatorname{Ob}Q$ and $\hom Q$ are the sets of elements and relations of $Q$, respectively. A poset $P$ is called a subposet of $Q$ if $\operatorname{Ob}P\subseteq\operatorname{Ob}Q$ and $\hom P\subseteq\hom Q$. In this case, we write $P\subseteq Q$. We call $P$ a full subposet if the inclusion functor $P\hookrightarrow Q$ is full, i.e., $p\leq p^{\prime}$ in $P$ whenever $p,p^{\prime}\in\operatorname{Ob}P$ and $p\leq p^{\prime}$ in $Q$.
Intersections of posets are defined in the obvious way, i.e., by taking intersections of both the underlying sets and the sets of relations. We define the union of posets analogously. In general a union of posets needn’t be a poset, as it may fail the transitivity property, but all unions considered in this paper satisfy transitivity.
The Hasse diagram of a finite poset $Q$ is the directed graph with vertex set $\operatorname{Ob}Q$ and an edge $(p,q)$ for every relation $p<q\in Q$ which does not factor as $p<r<q$ for some $r\in Q$.
Definition 2.2.
We call a category $C$ connected if $C$ is non-empty (i.e., has at least one object), and for each pair of objects $c,c^{\prime}\in\operatorname{Ob}C$, there exists a finite zigzag of morphisms
[…]
connecting $c$ and $c^{\prime}$. We call such a zigzag a path from $c$ to $c^{\prime}$. A component of $C$ is a maximal full, connected subcategory of $C$.
Note that a finite poset $Q$ is connected if and only if the undirected graph underlying its Hasse diagram is a connected graph.
Definition 2.3.
Given a poset $Z$, a nonempty subset $Q\subseteq Z$ is an interval if
  • $Q$ is connected,
  • whenever $p\leq q\leq r$ with $q\in Z$ and $p,r\in Q$, we have $q\in Q$.
We regard an interval $Q\subseteq Z$ as a full subposet of $Z$.
Definition 2.4.
Given a poset $Q$,
  • (i)
    the (closed) downset of $q\in Q$ is the set
    $\operatorname{dn}[q,Q]\coloneqq\{p\in Q\mid p\leq q\},$
  • (ii)
    the open downset of $q$ is
    $\operatorname{dn}(q,Q)\coloneqq\operatorname{dn}[q,Q]\setminus\{q\}=\{p\in Q\mid p<q\},$
  • (iii)
    For posets $P\subseteq Q$, the (closed) upset of $P$ in $Q$ is the set
    $\operatorname{up}[P,Q]\coloneqq\{q\in Q\mid q\geq p\textup{ for some }p\in P\}.$
Henceforth, we regard each of these as a full subposet of $Q$.
We let $M_{Q}$ denote the set of minima of $Q$.
When $Q$ is clear from context, we sometimes omit it from the above notation, i.e.,
$\operatorname{dn}(q)$ $\coloneqq\operatorname{dn}(q,Q),$
$\operatorname{dn}[q]$ $\coloneqq\operatorname{dn}[q,Q],$
$\operatorname{up}[P]$ $\coloneqq\operatorname{up}[P,Q],$
$M$ $\coloneqq M_{Q}.$
2.2. Limits
Given a small category $C$, a category $D$, and a functor $G\colon C\to D$, a cone on $G$ consists of an object $d\in\operatorname{Ob}D$, together with a morphism $\gamma_{c}\colon d\to G_{c}$ for each $c\in\operatorname{Ob}C$, called a cone map, such that for all morphisms $g\colon c\to c^{\prime}$ in $\hom C$, we have $\gamma_{c^{\prime}}=G_{g}\circ\gamma_{c}$. We write the cone as $(d,\gamma)$.
Definition 2.5.
The limit of $G$ is a cone $(d,\gamma)$ such that for any other cone $(d^{\prime},\gamma^{\prime})$, we have a unique morphism $\kappa\colon d^{\prime}\to d$ with $\gamma^{\prime}_{c}=\gamma_{c}\circ\kappa$ for all $c\in\operatorname{Ob}C$.
When they exist, limits are unique up to unique isomorphism, which justifies speaking of the limit of $G$. If $(d,\gamma)=\lim G$, we sometimes abuse terminology slightly and refer to $d$ as $\lim G$, particularly when $\gamma$ is clear from context. Cocones and colimits are defined dually.
Given functors $F\colon C\to D$ and $G\colon D\to E$ such that $\lim G$ and $\lim(G\circ F)$ both exist, the limit cone $(\lim G,\gamma)$ on $G$ restricts to a cone $(\lim G,\kappa)$ on $G\circ F$. The universality of $\lim(G\circ F)$ then yields a morphism $\lim G\to\lim(G\circ F)$ which commutes with the relevant cone maps. This is functorial in the sense that, given another functor $F^{\prime}\colon C^{\prime}\to C$, the morphism $\lim G\to\lim(G\circ F\circ F^{\prime})$ factors as
$\lim G\to\lim(G\circ F)\to\lim(G\circ F\circ F^{\prime}).$
The category $\mathbf{Vec}$ of vector spaces over a fixed field is complete and cocomplete, meaning that $\lim G$ and $\operatorname{colim}G$ exist for any $G\colon C\to\mathbf{Vec}$.
2.2.1. Limits as the Solutions to Linear Systems
For any small category $C$ and functor $G\colon C\to\mathbf{Vec}$, $\lim G$ admits a simple, concrete description as the solution of a system of linear equations: For $v\in\prod_{c\in\operatorname{Ob}C}G_{c}$ and $b\in\operatorname{Ob}C$, let $v_{b}$ be the projection of $v$ on $G_{b}$. For $\gamma\colon b\to c\in\hom C$, write $s(\gamma)=b$ and $t(\gamma)=c$. A standard result [mac2013categories, Theorem V.2.2] is that
(2.1) $\lim G=\left\{v\in\prod_{c\in\operatorname{Ob}C}G_{c}\ :\ G_{\gamma}(v_{s(\gamma)})=v_{t(\gamma)}\ \forall\,\gamma\in\hom C\right\},$
with each cone map $\lim G\to G_{c}$ the restriction of the projection
$\prod_{c\in\operatorname{Ob}C}G_{c}\twoheadrightarrow G_{c}.$
Elements of the vector space on the right side of Eq.˜2.1 are called sections of $G$.
In the special case of a functor $G\colon Q\to\mathbf{Vec}$, where $Q$ is a poset with finite downsets, we can express $\lim G$ as the solution to a smaller system of equations:
Proposition 2.6.
For a poset $Q$ with finite downsets and $G\colon Q\to\mathbf{Vec}$, we have
(2.2) $\lim G=\left\{v\in\!\prod_{m\in M}G_{m}\,:\,G_{mq}(v_{m})=G_{m^{\prime}q}(v_{m^{\prime}})\ \forall\,m,m^{\prime}\in M\textup{ with }m,m^{\prime}<q\right\},$
with each cone map $\lim G\to G_{q}$ given as the composition
$\lim G\hookrightarrow\prod_{m\in M}G_{m}\twoheadrightarrow G_{m^{\prime}}\xrightarrow{G_{m^{\prime}q}}G_{q},$
for any choice of $m^{\prime}\in M$ with $m^{\prime}\leq q$.
Definition 2.7.
  • (i)
    A presection of $G$ is an element of the vector space on the right side of Eq.˜2.2.
  • (ii)
    A presection basis of $G$ is a basis for this vector space.
Sketch of proof of Proposition˜2.6.
Each presection of $G$ extends uniquely to a section of $G$ via the structure maps of $G$. This defines an isomorphism from the vector space of presections to the vector space of sections. It is immediate that this isomorphism commutes with the cone maps. ∎
2.3. Initial Functors
Definition 2.8.
Given a functor $F\colon C\to D$ and $d\in\operatorname{Ob}D$, the comma category $(F\downarrow d)$ is the category whose
  • objects are the pairs $(c,\alpha)$, where $c\in\operatorname{Ob}C$ and $\alpha\in\hom(F(c),d)$,
  • morphisms $(c,\alpha)\to(c^{\prime},\alpha^{\prime})$ are morphisms $\kappa\colon c\to c^{\prime}$ in $C$ such that $\alpha=\alpha^{\prime}\circ F(\kappa)$.
Definition 2.9.
A functor $F\colon C\to D$ with $C,D$ small is initial if $(F\downarrow d)$ is connected for each $d\in\operatorname{Ob}D$.
The following standard result says that we can compute a limit by restricting a diagram along an initial functor.
Proposition 2.10 ([mac2013categories, Theorem IX.3.1]).
Consider functors $F\colon C\to D$ and $G\colon D\to E$ with $C$ and $D$ small and $F$ initial. If $\lim(G\circ F)$ exists, then $\lim G$ exists and the induced map $\lim G\to\lim(G\circ F)$ is an isomorphism in $E$.
Remark 2.11.
The proof of Proposition˜2.10 in [mac2013categories] gives a concrete construction of the cone $\lim G$ from the cone $\lim F\circ G$, where each cone map is taken to be the composition of a cone map of $\lim F\circ G$ with a structure map of $G$. This construction makes precise the idea that to compute $\lim G$, it suffices to compute $\lim G\circ F$.
We will be particularly interested in functors $F\colon C\to Q$ where $Q$ is a poset. In this case, for $q\in Q$, $(F\downarrow q)$ can be identified with the full subcategory of $C$ with objects $\{c\in\operatorname{Ob}C\mid F(c)\leq q\}$. In particular, we have the following:
Proposition 2.12.
An inclusion of posets $P\hookrightarrow Q$ is initial if and only if for each $q\in Q$, the intersection poset $\operatorname{dn}[q,Q]\cap P$ is connected.
Example 2.13.
Let $Q$ be the poset with underlying set
$Q=\{m_{1},m_{2},m_{3},m_{4},e_{1},e_{2},e_{3}\}$
whose Hasse diagram has edges
$\{(m_{1},e_{1}),(m_{2},e_{1}),(m_{2},e_{2}),(m_{3},e_{2}),(e_{1},e_{3}),(e_{2},e_{3}),(m_{4},e_{3})\}.$
For each $i\in\{1,2,3\}$, let $P_{i}\subseteq Q$ be the subposet with $\operatorname{Ob}P_{i}=\operatorname{Ob}Q$, whose Hasse diagram has edges
$\{(m_{1},e_{1}),(m_{2},e_{1}),(m_{2},e_{2}),(m_{3},e_{2}),(m_{i},e_{3}),(m_{4},e_{3})\};$
see Fig.˜1. Then each inclusion $P_{i}\hookrightarrow Q$ is an initial functor. Note that as posets $P_{1}\cong P_{3}\not\cong P_{2}$.
$m_{1}$$m_{2}$$m_{3}$$m_{4}$$e_{1}$$e_{2}$$e_{3}$
$m_{1}$$m_{2}$$m_{3}$$m_{4}$$e_{1}$$e_{2}$$e_{3}$
Figure 1. The Hasse diagram of the poset $Q$ from Example˜2.13 is shown on the left. The Hasse diagram of each subposet $P_{i}\subseteq Q$ has all of the blue edges and exactly one of the red edges shown on the right.
2.4. Functors from Posets to Vector Spaces
Recall that $\mathbf{Vec}$ denotes the category of vector spaces over a fixed field $\mathbbm{k}$. Given a poset $Q$, we often refer to a functor $G\colon Q\to\mathbf{Vec}$ as a $Q$-module. If $v\in G_{q}$, then we call $q$ the grade of $v$, and write $q=\mathrm{gr}(v)$. The $Q$-modules form an abelian category whose morphisms are the natural transformations.
A generating set of a $Q$-module $G$ is a set $S\subseteq\bigsqcup_{q\in Q}G_{q}$ such that for any $v\in\bigsqcup_{q\in Q}G_{q}$ we have
$v=\sum_{i=1}^{k}c_{i}G_{\mathrm{gr}(v_{i})\mathrm{gr}(v)}(v_{i})$
for some vectors $v_{1},v_{2},\dots,v_{k}\in S$ and scalars $c_{1},c_{2},\dots,c_{k}\in\mathbbm{k}$. The set $S$ is minimal if for all $v\in S$, the set $S\setminus\{v\}$ does not generate $G$. We say $G$ is finitely generated if there exists a finite generating set of $G$. Clearly, if $G$ is finitely generated, then a minimal generating set of $G$ exists.
Definition 2.14.
For $Z$ a poset and $Q\subseteq Z$ an interval, define the interval module $\mathbbm{k}^{Q}\colon Z\to\mathbf{Vec}$ by
$\mathbbm{k}^{Q}_{q}$ $=\begin{cases}\mathbbm{k}&\text{if }q\in Q\\ 0&\text{otherwise,}\end{cases}$ $\mathbbm{k}^{Q}_{p,q}=\begin{cases}\mathrm{Id}_{\mathbbm{k}}&{\text{if }p,q\in Q},\\ 0&{\text{otherwise}}.\end{cases}$
The interval module $\mathbbm{k}^{Q}$ is easily checked to be indecomposable.
2.4.1. Free Modules
A $Q$-module $G$ is free if there exists a multiset $\mathcal{B}$ of elements in $Q$ such that $G\cong\bigoplus_{q\in\mathcal{B}}\,\mathbbm{k}^{\operatorname{up}[\{q\},Q]}$. The rank of a free $Q$-module $G$ is $|\mathcal{B}|$, i.e., the number of indecomposable summands of $G$. A basis $B$ of a free $Q$-module $G$ is a minimal generating set. We define a function $\beta^{G}\colon Q\to\mathbb{N}$ by $\beta^{G}(q)=|\{s\in B\mid\mathrm{gr}(s)=q\}|$; an elementary linear algebra argument shows that $\beta^{G}$ is independent of the choice of $B$.
Given a morphism $\gamma\colon G\to G^{\prime}$ of finitely generated free $Q$-modules and a choice of ordered bases $B=\{b_{1},\ldots,b_{n}\}$ and $B^{\prime}=\{b^{\prime}_{1},\ldots,b^{\prime}_{m}\}$ for $G$ and $G^{\prime}$, we can represent $G$ as a matrix $\gamma]$, with each row and each column is labeled by an element of $Q$: The matrix is the representation of $\operatorname{colim}\gamma\colon\operatorname{colim}G\to\operatorname{colim}G^{\prime}$ with respect to the ordered bases induced by $B$ and $B^{\prime}$; the $i^{\mathrm{th}}$ column label is $\mathrm{gr}(b_{i})$; and the $i^{\mathrm{th}}$ row label is $\mathrm{gr}(b^{\prime}_{i})$. Regarding $\gamma$ as a functor $Q\times\{0,1\}\to\mathbf{Vec}$, one easily checks that $\gamma$ can be recovered up to isomorphism from $\gamma]$.
3. Main Definitions and Results
In this section we introduce the central definitions of this paper and state our main results, deferring the main proofs to subsequent sections.
3.1. Structure of Minimal Initial Functors
Definition 3.1.
An initial functor $F\colon C\to D$ is minimal if for any initial functor $F^{\prime}\colon C^{\prime}\to D$, we have
$|\operatorname{Ob}C|\leq|\operatorname{Ob}C^{\prime}|\quad\textup{and}\quad|\hom C|\leq|\hom C^{\prime}|.$
For $Q$ a poset, let
$I_{Q}=\{q\in Q\mid\operatorname{dn}(q)\textup{ is disconnected}\}.$
Our first main result, Theorem˜3.5 below, gives a simple characterization of the minimal initial functors $F\colon P\to Q$ for a large class of posets $Q$. Specifically, the theorem shows that such functors are the same, up to unique isomorphism, as initial scaffolds, which we now define.
Definition 3.2.
For $Q$ a poset with finite downsets, an initial scaffold of $Q$ is a subposet $j\colon P\hookrightarrow Q$ of the following form:
  • The set of elements of $P$ is $I_{Q}$.
  • For each $p\in P$ and each component $A$ of $\operatorname{dn}(p,Q)$, $P$ contains exactly one relation $m<p$ with $m\in A$. (Note that $m$ is a minimum of $Q$.)
By construction, all choices of an initial scaffold $P\subseteq Q$ have the same elements. However, the next example shows that $Q$ can have multiple non-isomorphic initial scaffolds.
Example 3.3.
In Example˜2.13, each of the three subposets $P_{1},P_{2},P_{3}\subseteq Q$ is an initial scaffold of $Q$.
Example 3.4.
It is easily checked that any interval $Q\subseteq\mathbb{N}^{2}$ has a unique initial scaffold $P$, which is the full subposet of $I_{Q}$ in $Q$. Further, $P$ is a zigzag poset, i.e., $P$ is isomorphic to a finite interval of the full subposet of $\mathbb{Z}^{2}$ with the following elements:
$\{(z,z)\mid z\in\mathbb{Z}\}\cup\{(z,z+1)\mid z\in\mathbb{Z}\}.$
Theorem 3.5 (Structure of Minimal Initial Functors).
For any poset $Q$ with finite downsets,
  • (i)
    an initial scaffold $j\colon P\hookrightarrow Q$ exists,
  • (ii)
    the inclusion $j$ is a minimal initial functor,
  • (iii)
    if $Q$ has a finite initial scaffold, then for any minimal initial functor $F\colon C\to Q$, there is an initial scaffold $j\colon P\hookrightarrow Q$ and a unique isomorphism $E\colon C\to P$ such that $F=j\circ E$:
[…]
We prove Theorem˜3.5 in Section˜4. Corollary˜3.11 below establishes that any interval $Q\subseteq\mathbb{N}^{d}$ has a finite initial scaffold. Thus, if $Q$ is finite or $Q\subseteq\mathbb{N}^{d}$ is an interval, then $Q$ satisfies the conditions of Theorem˜3.5.
Remark 3.6.
Theorem˜3.5 (ii) immediately implies a variant of the result for full subposets: If $Q$ is a poset with finite downsets and $P$ is the full subposet of $Q$ with $\operatorname{Ob}P=I_{Q}$, then the inclusion $P\hookrightarrow Q$ is initial. Moreover, for any full subposet $P^{\prime}\subseteq Q$ whose inclusion into $Q$ is initial, we have $P\subseteq P^{\prime}$. As illustrated by Example˜3.3, initial scaffolds needn’t be full, so this is a strictly weaker notion of minimality than that of Theorem˜3.5 (ii).
Remark 3.7 (Related work of Brüstle et al.).
In October 2025, Brüstle et al. posted a paper giving a variant of Remark˜3.6 under a weaker condition on the poset $Q$ [brustle2025generalized, Theorem C]. Previously, in October 2024, one of us, Lesnick, gave a talk on early versions of our results [Lesnick2024InitialHulls], which centered the result of Remark˜3.6. As one of the authors of [brustle2025generalized] was present for this talk, those authors attribute a version of their Theorem C to us in their paper.
In addition, a variant of our Proposition˜8.2 was obtained independently in [brustle2025generalized]. The paper [brustle2025generalized] also contains other related results which do not overlap with our work. In contrast to our work, [brustle2025generalized] does not deal explicitly with algorithmic questions.
Remark 3.8.
It is straightforward to check that for any poset $Q$ with finite downsets and initial scaffold $\subseteq Q, PP$ is the unique initial scaffold of itself. This provides an intrinsic characterization of posets arising as initial scaffolds.
3.2. Size of Initial Scaffolds
Recall from Section˜1 that we define $|P|$, the size of the initial scaffold $P\subseteq Q$, to be the number of relations in $P$, including identity relations. A natural way to study $|P|$ is to compare it to the number $n$ of minima of $Q$. The next example shows that $|P|$ can be arbitrarily large, even when $n=2$.
Example 3.9.
Let $Q=\{m_{1},m_{2},q_{1},\ldots,q_{k}\}$, where $m_{1}$ and $m_{2}$ are minima, $q_{1},\ldots,q_{k}$ are maxima, and $m_{i}\leq q_{j}$ for all $i$ and $j$. Then $Q$ is the unique initial scaffold of itself.
However, in the special case that $Q\subseteq\mathbb{N}^{d}$ is an interval, we show that $|P|$ is controlled by $m$, as follows:
Theorem 3.10.
Let $Q$ be an interval in $\mathbb{N}^{d}$ with $n$ minima and let $P\subseteq Q$ be an initial scaffold of $Q$. We have
  • (i)
    $|P|=\Theta(n)$ for $d\leq 3$,
  • (ii)
    $|P|=\Theta(n^{2})$ for $d>3$.
We prove Theorem˜3.10 in Section˜5. The case $d=3$ of Theorem˜3.10 is arguably the most interesting one. As noted in Section˜1, we prove the upper bounds of Theorem˜3.10 by applying an upper bound from [bayer1998monomial] on the size of monomial ideals. The lower bound of Theorem˜3.10 (i) is essentially trivial, while the lower bound of Theorem˜3.10 (ii) is obtained by an explicit construction.
Corollary 3.11.
Any interval $Q\subseteq\mathbb{N}^{d}$ has a finite initial scaffold.
Corollary˜3.11 is immediate from Theorem˜3.10 and the following standard result:
Lemma 3.12 (Dickson’s Lemma [dickson1913finiteness]).
Any full subposet of $\mathbb{N}^{d}$ has finitely many minima.
Proof.
This follows from the fact that polynomial rings are Noetherian. To elaborate, if $Q\subseteq\mathbb{N}^{d}$ is a full subposet, then $Q$ and $\operatorname{up}[Q,\mathbb{N}^{d}]$ have the same minima. The minima of $\operatorname{up}[Q,\mathbb{N}^{d}]$ are exactly the grades of a minimal set of generators of an ideal in the ring of polynomials $\mathbbm{k}[x_{1},\ldots,x_{d}]$; see Proposition˜5.10. Since this ring is Noetherian, every ideal is finitely generated. ∎
3.3. Computing Initial Scaffolds
We next state our bounds on the complexity of computing an initial scaffold, deferring the proofs and the algorithms underlying the bounds to Section˜6.
Our first bound is for an arbitrary finite poset:
Theorem 3.13.
Given the Hasse diagram $(V,E)$ of a finite poset $Q$, we can compute an initial scaffold of $Q$ in time $O(|V||E|)$.
The algorithm underlying Theorem˜3.13 is a straightforward application of depth-first search.
In the case that $Q\subseteq\mathbb{N}^{d}$ is a finite interval, every vertex is incident to $O(1)$ edges, so the bound of Theorem˜3.13 simplifies to $O(|V|^{2})=O(|\operatorname{Ob}Q|^{2})$. Since $|\operatorname{Ob}Q|$ may be quite large compared to the amount of data required to specify $Q$, one might hope to improve this bound. To this end, for $Q\subseteq\mathbb{N}^{d}$ an interval, we give two specialized algorithms for computing initial scaffolds. Both take as input a different (and usually smaller) representation of $Q$, which we now introduce:
Definition 3.14.
Given an interval $Q\subseteq\mathbb{N}^{d}$, let $Q^{\prime}\coloneqq\operatorname{up}[Q,\mathbb{N}^{d}]\setminus Q$. We call the pair $(M_{Q},M_{Q^{\prime}})$ of sets of minima the upset presentation of $Q$. We let $\|Q\|=|M_{Q}|+|M_{Q^{\prime}}|$.
It is easily checked that an interval $Q\subseteq\mathbb{N}^{d}$ is determined by its upset presentation. Informally, we think of $M_{Q}$ and $M_{Q^{\prime}}$ as the sets of birth and death points of $Q$, respectively. We emphasize that $\|Q\|$ and $|Q|$ are different quantities. Note that by Lemma˜3.12, $\|Q\|$ is always finite.
Example 3.15.
We give the upset presentations of three intervals in $\mathbb{N}^{2}$:
  • (i)
    If $Q=\{(0,0)\}$, then $M_{Q}=\{(0,0)\}$ and $M_{Q^{\prime}}=\{(1,0),(0,1)\}$.
  • (ii)
    If $Q=\mathbb{N}^{2}$, then $M_{Q}=\{(0,0)\}$ and $M_{Q^{\prime}}=\emptyset$.
  • (iii)
    If $Q=\{(z,0)\mid z\in\mathbb{N}\}$, then $M_{Q}=\{(0,0)\}$ and $M_{Q^{\prime}}=\{(0,1)\}$.
Remark 3.16.
If $Q$ is a finite interval in $\mathbb{N}^{d}$, then one can instead specify $Q$ by its extrema (i.e., sets of minimal and maximal elements). However, Example˜3.15 (ii) and (iii) illustrate that an infinite interval $Q\subseteq\mathbb{N}^{2}$ is generally not determined by its extrema.
Throughout, we assume that when a finite interval $Q\subset\mathbb{N}^{d}$ is specified by its extrema, the maxima and minima are given separately.
The following result bounds the cost of computing the initial scaffold of an interval in $\mathbb{N}^{d}$.
Theorem 3.17.
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=s$, we can compute an initial scaffold of $Q$ in time
  • (i)
    $O(s\log s)$ for $d=2,3$,
  • (ii)
    $O(s^{4})$ for $d>3$.
In the case that $Q$ is finite, these bounds also hold when $Q$ is instead specified by its extrema, with $s$ the total number of these.
Separate algorithms underlie the bounds (i) and (ii) of Theorem˜3.17. In the case $d=3$, the algorithm for (i) constructs the initial scaffold by iterating through two-dimensional slices of $\mathbb{N}^{3}$. The algorithm for (ii) hinges on the observation that it suffices to restrict attention to the subposet of $Q$ formed by minima and joins of pairs minima.
3.4. Computing Limits via Initial Scaffolds
We next consider the application of initial scaffolds to limit computation. First, we specify the precise definition of limit computation that we use in this paper:
Definition 3.18 (Limit Computation).
Let $Q$ be either a finite poset or an interval in $\mathbb{N}^{d}$, and let $G\colon Q\to\mathbf{Vec}$ be a functor. In this paper, computing $\lim G$ means computing a presection basis of $G$ (Definition˜2.7).
Remarks 3.19.
  • (i)
    While this definition of limit computation does not entail explicit computation of the cone maps of $\lim G$, Proposition˜2.6 makes clear that the cone maps are readily obtained from what we compute: Given a presection basis of $\lim G$, each cone map to a minimum is given by coordinate projection, while each cone map to a non-minimum is given by composing a coordinate projection with a structure map of $G$.
  • (ii)
    It is easily checked that if $j\colon P\hookrightarrow Q$ is an initial scaffold of $Q$, then the presection bases of $\lim G$ and $\lim(G\circ j)$ are identical. Therefore, according to Definition˜3.18, the problems of computing $\lim G$ and $\lim(G\circ j)$ are identical.
To give an algorithm for computing the limit of a functor $G\colon Q\to\mathbf{Vec}$, we must specify how such functors are represented in our computations. Here is one useful way:
Definition 3.20.
Given a poset $Q$ and $Q$-module $G$, a matrix representation of $G$ consists of:
  1. (1)
    a choice of ordered basis for each of the vector spaces $(G_{q})_{q\in Q}$ and
  2. (2)
    the matrices $G_{pq}]_{p\leq q\in Q}$ representing the structure maps $(G_{pq})_{p\leq q\in Q}$ with respect to these bases.
To bound the complexity of limit computation, we use the following well-known linear algebra result:
Proposition 3.21 ([jeannerod2013rank]).
A $y\times z$ matrix $A$ of rank $r$ can be transformed to row echelon in time $O(yzr^{\omega-2})$. Hence, using the standard $O(z^{2})$ backsolve algorithm, a basis for $\ker(A)$ can be computed in time $O(yz^{\omega-1})$.
We refer to the algorithm underlying Proposition˜3.21 as Gaussian Elimination, though it is different than the usual cubic-time version of Gaussian Elimination.
Together, Proposition˜3.21 and Proposition˜2.6 yield the following naive bound on the cost of limit computation, where
$r(G)\coloneqq\max_{q\in Q}\ \dim G_{q}.$
Proposition 3.22.
Let $Q$ be a finite poset with $n$ minima. Given a matrix representation of a $Q$-module $G$ with $r(G)=r$, $\lim G$ can be computed in time
$O(|Q|n^{\omega-1}r^{\omega}).$
Proof.
The solution to the system of equations in Proposition˜2.6 is the kernel of a linear system with $O(r|Q|)$ equations and $O(rn)$ variables. The result now follows from Proposition˜3.21. ∎
We can improve on Proposition˜3.22 by computing initial scaffolds. Specifically, given a poset $Q$ as in either Theorem˜3.13 or Theorem˜3.17 and $Q$-module $G$, we compute $\lim G$ in three steps:
  1. (1)
    Compute an initial scaffold $j\colon P\hookrightarrow Q$ using the algorithm of Theorem˜3.13 or Theorem˜3.17.
  2. (2)
    Compute a matrix representation of $G\circ j$.
  3. (3)
    Compute $\lim(G\circ j)$ by applying Proposition˜3.22.
Let $\Gamma$ denote the cost of the second step.
Corollary 3.23.
Given the Hasse diagram $(V,E)$ of a finite poset $Q$ with $n$ minima and a $Q$-module $G$ with $r(G)=r$, the above approach computes $\lim G$ in time
$O\left(|V||E|+|P|n^{\omega-1}r^{\omega}\right)+\Gamma.$
Proof.
By Theorem˜3.13, we can compute an initial scaffold of $Q$ in time $O(|V||E|)$. The result now follows from Proposition˜3.22. ∎
Corollary 3.24.
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=s$ and a $Q$-module $G$ with $r(G)=r$, the above approach computes $\lim G$ in time
  • (i)
    $O((rs)^{\omega})+\Gamma$ for $d=2,3$,
  • (ii)
    $O(r^{\omega}s^{\omega+1})+\Gamma$ for $d>3$.
Proof.
By Theorem˜3.17, we can compute an initial scaffold $j\colon P\hookrightarrow Q$ in time $O(s\log s)$ for $d=2,3$ and $O(s^{4})$ for $d>3$. In addition, Theorem˜3.10 tells us that $|P|=\Theta(s)$ if $d=2,3$ and $|P|=\Theta(s^{2})$ if $d>3$. The result now follows from Proposition˜3.22. ∎
The cost $\Gamma$ depends on the format in which the input diagram $G$ is given. For example, if we are given a matrix representation of $G$, then $\Gamma$ is negligible. However, in the TDA applications that initially motivated this work, we typically are not given matrix representation of $G$; instead, we are given a free presentation of $G$ or, more generally, a chain complex of free functors whose homology is isomorphic to $G$. In this case, we may apply Proposition˜3.26 below.
Definition 3.25.
For $Q$ a poset, a $(Q,r)$-complex is a chain complex
$X\xrightarrow{f}Y\xrightarrow{g}Z,$
of free $Q$-modules of total rank $r$, represented by labeled matrices $[f],[g]$ as in Section˜2.4.1. We call $\ker g/\operatorname{im}f$ the homology of the $(Q,r)$-complex.
Proposition 3.26.
Suppose we are given
  1. (1)
    a poset $Q$, represented in such a way that pairs of elements can be compared in constant time,
  2. (2)
    a $(Q,r)$-complex with homology $H$, and
  3. (3)
    the set of relations of a subposet $i\colon P\hookrightarrow Q$.
We can compute a matrix representation of $H\circ i$ in time $O(|P|r^{\omega})$.
We prove Proposition˜3.26 in Section˜7.1 via straightforward linear algebra.
Remark 3.27.
Note that if $g=0$, then the chain complex of Proposition˜3.26 is in fact a free presentation of $H$. In the general case where $g$ may be non-zero, one can in principle first compute a (minimal) presentation of $H$, and then take this as the input to algorithm underlying Proposition˜3.26. In TDA applications, it has been observed that the size of a minimal presentation of $H$ is often far smaller than the size of the original chain complex [fugacci2023compression].
In the case $Q=\mathbb{N}^{2}$ and in the case of $0^{\mathrm{th}}$ simplicial homology of filtrations indexed by arbitrary posets, efficient algorithms are known for computing a minimal presentation of $H$ [morozov2025computing, lesnick2022minimal, fugacci2023compression, bauer2023efficient]. In more general settings, the problem can be solved via Gröbner basis algorithms, e.g., Schreyer’s algorithm and its variants [schreyer1980berechnung, cox1998using, erocal2016refined, la1998strategies], but existing open-source implementations of such approaches are expensive on the types of large, sparse input considered in TDA [lesnick2022minimal].
Remark 3.28.
While some TDA constructions yield a chain complex whose chain modules are not free [botnan2023introduction, Section 5], such a chain complex can often readily be converted to a free chain complex with isomorphic homology [lesnick2015interactive, chacholski2017combinatorial].
Remark 3.29.
Given the Hasse diagram $(V,E)$ of a finite poset $Q$, we can compute a $|V|\times|V|$ binary matrix explicitly representing the partial order on $Q$ in time $O(|V||E|)$, by performing a breadth-first search starting from each $v\in V$. Using this matrix, we can check whether any relation $p\leq q$ holds in constant time, as required in Proposition˜3.26.
Together, Corollaries˜3.23, 3.13, 3.29 and 3.26 imply the following:
Corollary 3.30.
Given the Hasse diagram $(V,E)$ of a finite poset $Q$ with $n$ minima and a $(Q,r)$-complex with homology $H$, we can compute $\lim H$ in time
$O\left(|V||E|+|P|n^{\omega-1}r^{\omega}\right).$
Similarly, Corollaries˜3.24, 3.17 and 3.26 imply items (ii) and (iii) of the following corollary, which is an analogue of Corollary˜3.30 for intervals in $\mathbb{N}^{d}$.
Corollary 3.31.
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the upset presentation of an interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$ with $s=\|Q\|$, we can compute $\lim(H\circ i)$} in time
  • (i)
    $O(s\log s+r^{3})$ for $d=2$,
  • (ii)
    $O((rs)^{\omega})$ for $d=3$,
  • (iii)
    $O(r^{\omega}s^{\omega+1})$ for $d>3$.
We prove Corollary˜3.31 (i) in Section˜7.2 using a somewhat different approach than for parts (ii) and (iii): By Example˜3.4, an interval in $\mathbb{N}^{2}$ has a unique initial scaffold, which is a zigzag poset. This allows us to compute the limit using zigzag persistence computation, rather than the general limit computation of Proposition˜3.22, and this turns out to be more efficient.
Remark 3.32.
Corollaries˜3.31 and 3.24 also hold if $Q$ is instead specified by its minima and maxima and $s$ is the total number of these.
Remark 3.33.
All of the above definitions and results dualize immediately, in particular, yielding a final scaffold, which is a minimal final functor that can be used to compute colimits.
3.5. Generalized Rank Computation
As an application of our main results, we consider the problem of computing the generalized rank of a diagram of vector spaces, a problem arising in TDA which has previously been studied in [DKM22, dey2024computing, asashiba2024interval].
Definition 3.34 ([kim2018generalized]).
For a connected poset $Q$ and a functor $G\colon Q\to\mathbf{Vec}$, the composition of cone and cocone maps $\lim G\to G_{q}\to\operatorname{colim}G$ is independent of the choice of $q\in Q$. The rank of this map, which we denote $\operatorname{grank}(G)$, is called the generalized rank of $G$.
Generalized ranks are of interest in TDA for three reasons.
  1. (1)
    The invariant $\operatorname{grank}(G)$ has a simple representation-theoretic interpretation: It is the number of interval summands of $G$ with support $Q$ [chambers2018persistent, Corollary 3.2].
  2. (2)
    Recent work by Xin et al. [xin2023gril] has used the generalized ranks along certain subintervals of $\mathbb{N}^{2}$ called worms to define GRIL, a novel vectorization of 2-parameter persistence for supervised learning.
  3. (3)
    Applying Möbius inversion to the function sending an interval $j\colon Z\hookrightarrow Q$ to $\operatorname{grank}(G\circ j)$ yields an invariant of $G$ called the generalized persistence diagram [kim2018generalized], a generalization of the usual persistence diagram of a $\mathbb{N}$-module that has been actively studied in recent work [botnan2024signed, kim2024bigraded, kim2025super, carriere2025sparsification, asashiba2023approximation]. The large size of generalized persistence diagrams is an obstacle to their practical use [kim2025super], but they are of interest from a theoretical standpoint.
Before stating our results on generalized rank computation, we briefly discuss prior work on this. Dey, Kim, and Mémoli [DKM22] gave the following bound on the cost of computing a generalized rank over an interval in $\mathbb{N}^{2}$:
Theorem 3.35 ([DKM22]).
Given an $(\mathbb{N}^{2},r)$-complex with homology $H$ and the $s$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{2}$, we can compute $\operatorname{grank}(H\circ i)$ in time $O((s+r)^{\omega})$.
The algorithm underlying Theorem˜3.35 determines $\operatorname{grank}(G)$ by computing the zigzag barcode of $G\circ j$, where $j\colon P\hookrightarrow Q$ is a zigzag containing the initial and final scaffolds of $Q$. It is shown in [DKM22] that $\operatorname{grank}(G)$ is the number of copies of the full interval $P$ in this barcode.
Inspired by Theorem˜3.35, Dey and Xin [dey2024computing] gave an algorithm to compute the generalized rank over a general poset by reducing the problem to a zigzag persistence computation. However, it is actually asymptotically more efficient to compute the generalized rank by directly computing the limits and colimits, even without the speedups enabled by initial and final functors.
For $Q$ a finite poset, $i\colon P\hookrightarrow Q$ an interval, and $G$ a $Q$-module, Asashiba and Liu [asashiba2024interval] gave an explicit formula for $\operatorname{grank}(G\circ i)$ as the difference of the rank of two matrices. In that work, the formalism of (co)limits is not explicitly used, and a bound on the cost of computing $\operatorname{grank}(G\circ i)$ is not explicitly given.
By computing initial and final scaffolds, we obtain the following bounds, whose proofs we give in Section˜8:
Corollary 3.36.
Let $Q$ be a finite, connected poset with $s$ extrema and let $^{\mathcal{I}}, PP^{\mathcal{F}}$ denote initial and final scaffolds of $Q$. Given the Hasse diagram $(V,E)$ of $Q$ and a $(Q,r)$-complex with homology $H$, we can compute $\operatorname{grank}(H)$ in time
$O(|V||E|+(|P^{\mathcal{I}}|+|P^{\mathcal{F}}|)s^{\omega-1}r^{\omega}).$
Corollary 3.37.
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the $s$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$, we can compute $\operatorname{grank}(H\circ i)$ in time
  • (i)
    $O(s\log s+r^{\omega})$ for $d=2$,
  • (ii)
    $O((rs)^{\omega})$ for $d=3$,
  • (iii)
    $O(r^{\omega}s^{\omega+1})$ for $d>3$.
Note that Corollary˜3.37 (i) slightly strengthens Theorem˜3.35. As we explain in Section˜8, Corollary˜3.36 and Corollary˜3.37 (ii) and (iii) follow readily from our main results, via straightforward linear algebra. To prove Corollary˜3.37 (i), we use an additional argument to cast the generalized rank computation as a zigzag persistence computation. This argument is similar to the one in [DKM22], though we frame it in the general formalism of initial and final scaffolds.
4. Proof of Structure Theorem for Minimal Initial Functors
In this section, we prove Theorem˜3.5. We prepare for the proof with several definitions and preliminary results.
Let $Q$ be a poset with finite downsets. In this section, all open and closed downsets are taken with respect to $Q$, so we suppress $Q$ in our notation for downsets. Recall from Definition˜3.2 that for any initial scaffold $P\subseteq Q$, we have $\operatorname{Ob}P=I_{Q}$. Since
$M_{Q}=\{q\in Q\mid\operatorname{dn}(q)\textup{ is empty}\}$
and the empty set is disconnected (see Definition˜2.2), we have $M_{Q}\subseteq I_{Q}$. Let $E_{Q}=I_{Q}\setminus M_{Q}$, i.e.,
$E_{Q}=\{q\in Q\mid\operatorname{dn}(q)\textup{ is disconnected and non-empty}\}.$
We call elements of $E_{Q}$ essential.
Definition 4.1.
Given a functor $F\colon C\to D$, the image of $F$ is the pair
$\operatorname{im}(F)\coloneqq(F(\operatorname{Ob}C),F(\hom C)),$
where $F(\operatorname{Ob}C)\subseteq\operatorname{Ob}D$ and $F(\hom C)\subseteq\hom D$.
We note that $\operatorname{im}F$ is not necessarily a subcategory of $D$, since $F(\hom C)$ may not be closed under composition; see Example˜4.4 below. We sometimes abuse notation slightly and, e.g., write $d\in\operatorname{im}F$ to mean $d\in F(\operatorname{Ob}C)$, or $\gamma\in\operatorname{im}F$ to mean $\gamma\in F(\hom C)$.
Since the definition of a connected category (Definition˜2.2) does not involve composition of morphisms, it extends to any collections of objects and morphisms in an ambient category. In particular, the notion of connectivity makes sense for $\operatorname{im}F$. We use this to state the next proposition, whose straightforward proof we omit.
Proposition 4.2.
If a category $C$ is connected and $F\colon C\to D$ is any functor, then $\operatorname{im}(F)$ is also connected.
We let $\overline{\operatorname{im}F}$ denote the subcategory of $D$ generated by $\operatorname{im}F$, i.e., the category obtained from $\operatorname{im}F$ by including the compositions of all finite chains of composable morphisms in $\operatorname{im}F$.
Proposition 4.3.
If a poset $Q$ has finite downsets, then for any initial functor $F\colon C\to Q$,
  • (i)
    $I_{Q}\subseteq{\operatorname{im}F}$,
  • (ii)
    there exists an initial scaffold $P\subseteq Q$ with $P\subseteq\overline{\operatorname{im}F}$.
Proof.
To prove (i), first recall from Section˜2.3 that for any functor $F\colon C\to Q$ and $q\in Q$, $(F\downarrow q)$ can be identified with the full subcategory of $C$ with objects $\{c\in\operatorname{Ob}c\mid F(c)\leq q\}$. If $F$ is initial, then by definition $(F\downarrow q)$ is connected, hence non-empty, so $M_{Q}\subseteq\operatorname{im}(F)$.
Next we show that $E_{Q}\subseteq\operatorname{im}F$. To do so, we assume to the contrary that there exists $q\in E_{Q}$ with $q\not\in\operatorname{im}F$, and show that then $(F\downarrow q)$ is disconnected, a contradiction. Let $\hat{F}\colon(F\downarrow q)\to Q$} be the restriction of $F$ to $(F\downarrow q)$ and note that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}[q]$. Since we assume $q\not\in\operatorname{im}F$, we have $\operatorname{im}F\cap\operatorname{dn}[q]\subseteq\operatorname{dn}(q)$. Moreover, $\operatorname{im}F$ contains all minima of $Q$, so since $Q$ has finite downsets, $\operatorname{im}F$ has non-empty intersection with each component of $\operatorname{dn}(q)$. Since $q$ is essential, $\operatorname{dn}(q)$ is disconnected. It follows that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}(q)$ is disconnected. Thus, $(F\downarrow q)$ is disconnected by the contrapositive of Proposition˜4.2, a contradiction. We conclude that $E_{Q}\subseteq\operatorname{im}F$, and hence that $I_{Q}\subseteq F(\operatorname{Ob}C)$, establishing (i).
In view of (i), proving (ii) amounts to showing that for each $q\in E_{Q}$ and component $A$ of $\operatorname{dn}(q)$, there is a relation $m<q$ in $\overline{\operatorname{im}F}$ with $m$ a minimum in $A$. Since $Q$ has finite downsets and $M_{Q}\subseteq I_{Q}\subseteq\operatorname{im}F$ by (i), we have $\operatorname{im}F\cap A\neq\emptyset$. Moreover, since $(F\downarrow q)$ is connected, Proposition˜4.2 implies that $\operatorname{im}\hat{F}=\operatorname{im}F\cap\operatorname{dn}[q]$ is connected. Thus, since $q\in\operatorname{im}F$, there exists at least one relation $a_{0}<q$ in $\operatorname{im}F$ where $a_{0}\in A$. If $a_{0}\in M_{Q}$, we are done. Otherwise, by applying the same argument to any component of $\operatorname{dn}(a_{0})$, we obtain a relation $a_{1}<a_{0}$ in $\operatorname{im}F$ with $a_{1}\in A$; indeed to show the existence of $a_{0}$, we did not use the fact that $q$ is essential, so the same argument applies. Since $Q$ has finite downsets, by continuing in this way, we eventually obtain a string of relations
$m<\cdots<a_{1}<a_{0}<q$
in $\operatorname{im}F$ with $m\in M_{Q}\cap A$. We then have $m<q$ in $\overline{\operatorname{im}F}$ as desired. ∎
The following example shows that Proposition˜4.3 (ii) is no longer true if we replace $\overline{\operatorname{im}F}$ with $\operatorname{im}F$ in the statement.
Example 4.4.
Consider the posets $R=\{a,b,c,d,e\}$ and $Q=\{w,x,y,z\}$ whose Hasse diagrams are as shown in Fig.˜2.
R:R:aabbccddeeQ:Q:wwxxyyzz
Figure 2. The initial functor $F\colon R\to Q$ of Example˜4.4 maps each element of $R$ to the element of $Q$ directly below it. For example, $F(b)=F(c)=x$.
$Q$ has a unique initial scaffold $P\subseteq Q$, which is the full subposet with objects $\{w,y,z\}$. Let $F\colon R\to Q$ be the functor given by
F(a)=w,F(b)=F(c)=x,F(d)=y,F(e)=x.F(a)=w,\ F(b)=F(c)=x,\ F(d)=y,\ F(e)=x.
$F$ is easily checked to be initial, but $\operatorname{im}(F)$ does not contain the relation $w\leq y$, which is in $P$.
Given any category $C$, there is an associated thin category $\operatorname{thin}(C)$ with the same objects such that $\hom(c,c^{\prime})$ is empty in $\operatorname{thin}(C)$ if and only if $\hom(c,c^{\prime})$ is empty in $C$. Moreover, we have a unique functor $\pi\colon C\to\operatorname{thin}(C)$ which is the identity on objects.
If $D$ is a thin category (say a poset), then any functor $F\colon C\to D$ induces a functor $\operatorname{thin}(F)\colon\operatorname{thin}(C)\to D$ satisfying $F=\operatorname{thin}(F)\circ\pi$.
Lemma 4.5.
If $F\colon C\to D$ is initial and $D$ is thin, then $\operatorname{thin}(F)$ is initial.
Proof.
For any $d\in D$, the categories $(F\downarrow d)$ and $(\operatorname{thin}(F)\downarrow d)$ have the same objects, since $C$ and $\operatorname{thin}(C)$ have the same objects, while $F$ and $\operatorname{thin}(F)$ act identically on objects. In particular, since $(F\downarrow d)$ is nonempty, so is $(\operatorname{thin}(F)\downarrow d)$. To show that $(\operatorname{thin}(F)\downarrow d)$ is connected, consider $(c,\gamma),(c^{\prime},\gamma^{\prime})$ in $(\operatorname{thin}(F)\downarrow d)$. These are also objects in the connected category $(F\downarrow d)$, so there is a path in $(F\downarrow d)$ connecting $(c,\gamma)$ and $(c^{\prime},\gamma^{\prime})$. Applying $\pi$ yields a path in $(\operatorname{thin}(F)\downarrow d)$ connecting $(c,\gamma)$ and $(c^{\prime},\gamma^{\prime})$. Hence $(\operatorname{thin}(F)\downarrow d)$ is connected. ∎
Lemma 4.6.
Given a poset $A$, let $\mathcal{W}$ be a set of connected subposets of $A$ with $A=\bigcup_{P\in\mathcal{W}}P$. Then $A$ is connected if and only if for each $x,y\in A$, there exists a sequence of posets $P_{1},\ldots P_{l}$ in $\mathcal{W}$ such that $x\in P_{1}$, $y\in P_{l}$, and each $P_{i}\cap P_{i+1}$ is non-empty.
Proof.
If $A$ is connected and $x,y\in A$, then there is a path in $A$ of the form
[…]
Let $r_{i}$ denote the $i^{\mathrm{th}}$ relation in this path, starting from the left. By the way $\mathcal{W}$ is defined, $r_{i}$ belongs to some $P_{i}\in\mathcal{W}$. Then the sequence $P_{1},\ldots,P_{l}$ has the required properties. Conversely, assume that for each $x,y\in A$ there exists a sequence of $P_{1},\ldots,P_{l}$ of subposets in $\mathcal{W}$ as in the theorem statement. Choose a sequence of elements $x=a_{0},a_{1},\ldots,a_{l}=y$, such that $a_{i}\in P_{i}\cap P_{i+1}$ for $i\in\{1,\ldots,l-1\}$. Then for each $i\in\{1,\ldots,l\}$, we have $a_{i-1},a_{i}\in P_{i}$ so since $P_{i}$ is connected, there exists a path in $P_{i}$ from $a_{i-1}$ to $a_{i}$. These paths assemble into a path in $A$ from $x$ to $y$. Hence $A$ is connected. ∎
We are now ready to prove Theorem˜3.5. We first recall the statement:
Theorem 3.5.
For any poset $Q$ with finite downsets,
  • (i)
    an initial scaffold $j\colon P\hookrightarrow Q$ exists,
  • (ii)
    the inclusion $j$ is a minimal initial functor,
  • (iii)
    if $Q$ has a finite initial scaffold, then for any minimal initial functor $F\colon C\to Q$, there is an initial scaffold $j\colon P\hookrightarrow Q$ and a unique isomorphism $E\colon C\to P$ such that $F=j\circ E$.
Proof of Theorem˜3.5.
Since $Q$ has finite downsets, for all $q\in Q$, each component $A$ of $\operatorname{dn}(q)$ contains a minimum. Hence an initial scaffold $j\colon P\to Q$ always exists; this gives (i). Alternatively, (i) follows from Proposition˜4.3 (ii), since there always exists an initial functor with target $Q$, e.g., the identity functor.
In view of Proposition˜2.12, showing that $j$ is initial amounts to checking that for each $q\in Q$, the intersection $\operatorname{dn}[q]\cap P$ is connected. We proceed by induction on $q$ with respect to the partial order on $\operatorname{dn}[q]$; such an induction makes sense because $Q$ has finite downsets. The base case is that $q$ is a minimum. Then $q\in P$, so
$\operatorname{dn}[q]=\{q\}=\operatorname{dn}[q]\cap P$
is connected. For the induction step, consider $q\in Q$ which is not a minimum and assume that $\operatorname{dn}[q^{\prime}]\cap P$ is connected for all $q^{\prime}<q$. Let
$A_{1},\ldots,A_{k}$
be the connected components of $\operatorname{dn}(q)$.
We claim that for each $i$, $A_{i}\cap P$ is connected. Since $Q$ has finite downsets, there is a finite set $S\subseteq A_{i}$ such that
$A_{i}=\bigcup_{s\in S}\operatorname{dn}[s].$
Since $A_{i}$ is connected, as is each downset $\operatorname{dn}[s]$, Lemma˜4.6 implies that for each $x,y\in A_{i}\cap P$ there is a sequence $s_{1},\ldots,s_{l}$ of elements in $S$ such that $x\in\operatorname{dn}[s_{1}]$, $y\in\operatorname{dn}[s_{l}]$, and
$\operatorname{dn}[s_{i}]\cap\operatorname{dn}[s_{i+1}]$
is non-empty for each $i$. Each such intersection contains a minimum of $Q$ and hence an element of $P$, since $P$ contains all minima. Thus, letting $P_{i}=\operatorname{dn}[s_{i}]\cap P$, each intersection $P_{i}\cap P_{i+1}$ is non-empty. By the induction hypothesis, each $P_{i}$ is connected. Noting that
$A_{i}\cap P=\left(\bigcup_{s\in S}\operatorname{dn}[s]\right)\cap P=\bigcup_{s\in S}\operatorname{dn}[s]\cap P,$
applying Lemma˜4.6 with $A=A_{i}\cap P$ and $\mathcal{W}=\{\operatorname{dn}[s]\cap P\}_{s\in S}$ gives that $A_{i}\cap P$ is connected, as claimed.
If $q\not\in P$, then we have $k=1$ and so
$\operatorname{dn}[q]\cap P=\operatorname{dn}(q)\cap P=A_{1}\cap P$
is connected. If $q\in P$, then $q$ is essential. Thus, by the definition of an initial scaffold, for each $i$ we have a relation $m_{i}<q$ in $P$ with $m_{i}\in A_{i}$. Moreover, $\operatorname{dn}[q]\cap P$ is the poset obtained from
$\operatorname{dn}(q)\cap P=\bigcup_{i=1}^{k}(A_{i}\cap P)$
by inserting $q$, together with the relations $m_{i}<q$ for each $i$. Since each $A_{i}\cap P$ is connected, it follows that $\operatorname{dn}[q]\cap P$ is connected. This finishes the proof that $j$ is initial.
To see that $j$ is minimal, note that for any initial functor $F\colon C\to Q$, Proposition˜4.3 (i) implies that $|\operatorname{Ob}P|=|I_{Q}|\leq|\operatorname{Ob}C|$. Moreover, Proposition˜4.3 (ii) implies that for every $q\in E_{Q}$ and component $A$ of $\operatorname{dn}(q)$ there is a relation $a<q$ in $\operatorname{im}F$ with $a\in A$. It follows that $|\hom P|\leq|\hom C|$. Hence $j$ is minimal. This proves (ii).
To prove (iii), assume that $Q$ has a finite initial scaffold (i.e., $I_{Q}$ is finite) and let $F\colon C\to Q$ be a minimal initial functor. Then (ii) and Proposition˜4.3 (i) together imply that $F$ is an injection on objects. In addition, Lemma˜4.5 implies that $C$ is thin, since if $C$ is not already thin, then $|\hom(\operatorname{thin}(C))|<|\hom C|$, contradicting that $F$ is minimal. Therefore, $F$ is injective on both objects and morphisms, so $\operatorname{im}F$ is a subcategory of $Q$, and in particular, is closed under composition. That is, $\overline{\operatorname{im}F}=\operatorname{im}F$. By Proposition˜4.3 (ii), $\overline{\operatorname{im}F}=\operatorname{im}F$ contains an initial scaffold $P$. The minimality of $F$ now implies that $\operatorname{im}(F)=P$. It follows that we have a unique isomorphism $E\colon C\to P$ satisfying $F=j\circ E$; namely, $E$ is the functor obtained from $F$ by restricting the target to $\operatorname{im}F=P$. ∎
In the induction underlying our proof of Theorem˜3.5 (ii), we have also essentially proven the following result, which will be useful in our algorithm to compute initial scaffolds of intervals in $\mathbb{N}^{d}$; see Section˜6.3.
Proposition 4.7.
Given a poset $Q$ with finite downsets, an initial scaffold $P\subseteq Q$, and $q\in Q$, the inclusion $P\cap\operatorname{dn}(q)\hookrightarrow\operatorname{dn}(q)$ induces a bijection on path components.
Proof.
If $q$ is a minimum, then $\operatorname{dn}(q)$ is empty, so this is clear. If $q$ is not a minimum, the statement is equivalent to the statement that for each connected component $A_{i}$ of $\operatorname{dn}(q)$, $P\cap A_{i}$ is connected. We have shown this in the inductive step of the proof that $P\hookrightarrow Q$ is initial (Theorem˜3.5 (ii)). ∎
5. Size of Initial Scaffolds of Intervals in $\mathbb{N}^{d}$
In this section, we prove Theorem˜3.10. Recall that this theorem gives tight bounds on the size of an initial scaffold of an interval $Q\subseteq\mathbb{\mathbb{N}}^{d}$, in terms of the number of minima of $Q$. The proof uses well-known results about the Betti numbers of monomial ideals. We begin by reviewing these.
5.1. Betti Numbers of Monomial Ideals
Given a poset $Q$, a free resolution $X$ of a $Q$-module $G$ is an exact sequence of free $Q$-modules
$X=\quad\cdots\xrightarrow{\partial_{3}}X_{2}\xrightarrow{\partial_{2}}X_{1}\xrightarrow{\partial_{1}}X_{0}$
such that $G\cong\mathrm{coker}(\partial_{1})$. $X$ is minimal if any free resolution of $G$ has a direct summand isomorphic to $X$.
A standard structure theorem [peeva2010graded, Sections 1.7 and 1.9] says that if $G\colon\mathbb{N}^{d}\to\mathbf{Vec}$ is finitely generated, then there exists a minimal resolution $X$ of $G$ which is unique up to isomorphism, and each $X_{i}$ is finitely generated. Moreover, Hilbert’s Syzygy Theorem says that $X_{i}=0$ for $i>d$. Recalling the notation of Section˜2.4.1, we call the function $\beta^{X_{i}}\colon\mathbb{N}^{d}\to\mathbb{N}$ the $i^{\mathrm{th}}$ (multigraded) Betti number of $G$, and denote it as $\beta_{i}^{G}$.
Definition 5.1.
Let $R^{d}=\mathbbm{k}[x_{1},\dots,x_{d}]$ denote the polynomial ring in $d$ variables. A monomial is an element of $R^{d}$ of the form
$x^{z}\coloneqq x_{1}^{z_{1}}x_{2}^{z_{2}}\cdots x_{d}^{z_{d}}$
for some $z=(z_{1},\ldots,z_{d})\in\mathbb{N}^{d}$. An ideal $J\subseteq R^{d}$ generated by a set of monomials is called a monomial ideal.
We regard a monomial ideal $J$ as an $\mathbb{N}^{d}$-indexed persistence module with
$J_{z}=\begin{cases}\mathbbm{k}x^{z}&\textup{ if }x^{z}\in J,\\ 0&\textup{ otherwise},\end{cases}$
and each structure map $J_{yz}\colon J_{y}\to J_{z}$ given by multiplication with the monomial $x^{z-y}$.
A version of a well-known result called Hochster’s formula [miller2005combinatorial, Theorem 1.34] gives the Betti numbers of a monomial ideal $J$ in terms of the reduced homology of a simplicial complex built from $J$. To state the result, we need the following notation and definition: Given $s\subseteq\{1,...,d\}$, let $\mathbf{e}_{s}=\sum_{j\in s}\mathbf{e}_{j}\subseteq\{0,1\}^{d}$, where $\mathbf{e}_{j}\subseteq\mathbb{N}^{d}$ denotes the $j^{\mathrm{th}}$ standard basis vector.
Definition 5.2.
For $J\subseteq R^{d}$ an ideal and $z\in\mathbb{N}^{d}$, define the Upper Koszul complex of $J$ at $z$ to be the simplicial complex
$K^{z}(J)=\{s\subseteq\{1,\ldots,d\}\mid s\neq\emptyset,\ x^{z-\mathbf{e}_{s}}\in J\}.$
Thus, vertices in $K^{z}(J)$ are indices $j\subseteq\{1,\ldots,d\}$ such that $x^{z-\mathbf{e}_{j}}\in J$ and edges are pairs of coordinates $\{j,k\}$ such that $x^{z-\mathbf{e}_{j}-\mathbf{e}_{k}}\in J$. In the following, $\tilde{H}$ denotes reduced simplicial homology with coefficients in $\mathbbm{k}$.
Theorem 5.3 ([miller2005combinatorial, Theorem 1.34]).
For any monomial ideal $J$, $i\geq 0$, and $z\in\mathbb{N}^{d}$, we have $\beta_{i}^{J}(z)=\dim\tilde{H}_{i-1}(K^{z}(J))$.
Let $\pi_{0}(W)$ denote the set of connected components of a simplicial complex $W$. In the case $i=1$, standard properties of reduced homology yield the following corollary of Theorem˜5.3:
Corollary 5.4.
If $K^{z}(J)$ is non-empty, then
$\beta_{1}^{J}(z)=|\pi_{0}(K^{z}(J))|-1,$
and if $K^{z}(J)$ is empty, then $\beta_{1}^{J}(z)=0$.
5.2. Bounds on Betti Numbers of Monomial Ideals
A well-known result of Bayer, Peeva, and Sturmfels [bayer1998monomial] provides bounds on the Betti numbers of monomial ideals. See also [miller2005combinatorial] for an expository treatment. We state an asymptotic version of the result, leveraging a standard upper bound on the number of $i$-dimensional faces of a convex polytope [edelsbrunner1987algorithms, Theorem 6.12].
Theorem 5.5 ([bayer1998monomial]).
Suppose $J\subseteq R^{d}$ is a monomial ideal with a minimal generating set of size $n$. Then for each $i$,
$\sum_{z\in\mathbb{N}^{d}}\beta^{J}_{i}(z)=O(n^{\min(i,\lfloor\frac{d}{2}\rfloor)}).$
Remark 5.6.
Theorem˜5.5 is proven by showing that $\sum_{z\in\mathbb{N}^{d}}\beta^{J}_{i}(z)$ is at most the number of $i$-dimensional faces of a polytope in $\mathbb{R}^{d}$ with $n$ vertices. McMullen’s famous upper bound theorem [mcmullen1970maximum], which asserts that the maximum number of $i$-dimensional faces is attained by a cyclic polytope, gives a tight, non-asymptotic bound on the number of such faces. This leads to a non-asymptotic version of Theorem˜5.5, which appears in [bayer1998monomial, miller2005combinatorial]. By instead using the asymptotic bound on the number of $i$-dimensional faces given in [edelsbrunner1987algorithms, Theorem 6.12], we obtain Theorem˜5.5.
We do not know if Theorem˜5.5 is tight in general, but we observe that it is tight in the case $i=1$:
Corollary 5.7.
For $J$ and $n$ as in Theorem˜5.5, we have
$\sum_{z\in\mathbb{N}^{d}}\beta^{J}_{1}(z)=\begin{cases}\Theta(n)&\textup{ for }d\leq 3,\\ \Theta(n^{2})&\textup{ for }d>3.\end{cases}$
Proof.
Theorem˜5.5 implies that
$\sum_{z\in\mathbb{N}^{d}}\beta^{J}_{1}(z)=\begin{cases}O(n)&\textup{ for }d\leq 3,\\ O(n^{2})&\textup{ for }d>3.\end{cases}$
In the case $d=1$, we always have $n=1$, so the bound $\Theta(n)$ holds trivially. It is easily checked that in the case $d=2$, we have
$\sum_{z\in\mathbb{N}^{2}}\beta^{J}_{1}(z)=n-1.$
Thus, the bound $\Theta(n)$ holds for $d=2$. Example˜5.14 below gives a family of monomial ideals in $\mathbb{N}^{4}$ for which the sum of first Betti numbers grows quadratically with the number of minima. Thus, the bound $\Theta(n^{2})$ holds for $d=4$. Finally, for any monomial ideal $J\subseteq R^{d}$ and $d^{\prime}>d$, note that $J^{\prime}\coloneqq J\times R^{d^{\prime}-d}$ is an ideal of $R^{d^{\prime}}$ such that for all $i$,
$\sum_{z\in\mathbb{N}^{d}}\beta^{J}_{i}(z)=\sum_{z\in\mathbb{N}^{d^{\prime}}}\beta^{J^{\prime}}_{i}(z).$
Thus, the claimed bounds hold for all $d$. ∎
5.3. Size Bounds
Our proof of Theorem˜3.10 will make essential use of the following result, which will also be useful in Sections˜6.3 and 6.2, when we consider the problem of computing an initial scaffold of an interval in $\mathbb{N}^{d}$.
All upsets below are taken in the poset $\mathbb{N}^{d}$.
Lemma 5.8.
For $Q\subseteq\mathbb{N}^{d}$ an interval, $I_{Q}=I_{\operatorname{up}[Q]}\cap Q$. In particular, $I_{Q}\subseteq I_{\operatorname{up}[Q]}$.
Proof.
It is clear that $M_{Q}=M_{\operatorname{up}[Q]}=M_{\operatorname{up}[Q]}\cap Q$. Therefore, it suffices to check that $E_{Q}=E_{\operatorname{up}[Q]}\cap Q$.
We claim that for all $q\in Q$, $\operatorname{dn}(q,Q)=\operatorname{dn}(q,\operatorname{up}[Q])$. To check the claim, first note that $\operatorname{dn}(q,Q)\subseteq\operatorname{dn}(q,\operatorname{up}[Q])$ because $Q\subseteq\operatorname{up}[Q]$. Conversely, suppose $p\in\operatorname{dn}(q,\operatorname{up}[Q])$. Since $\mathbb{N}^{d}$ is bounded below, there exists $m\in M_{\operatorname{up}[Q]}=M_{Q}$ with $m\leq p<q$. Since $Q$ is an interval, we have $p\in Q$. Thus, $\operatorname{dn}(q,\operatorname{up}[Q])\subseteq\operatorname{dn}(q,Q)$ which establishes the claim.
The claim implies that for $q\in Q$, we have $q\in E_{Q}$ if and only if $q\in E_{\operatorname{up}[Q]}$. Thus, $E_{Q}=E_{\operatorname{up}[Q]}\cap Q$, as desired. ∎
It can be shown that the containment $I_{Q}\subseteq I_{\operatorname{up}[Q]}$ of Lemma˜5.8 is in fact an equality when $d\leq 2$, though we do not need this. The next example shows that this containment can be strict when $d=3$.
Example 5.9.
Consider the following interval in $\mathbb{N}^{3}$,
$Q=\{(1,0,0),(0,1,0),(0,0,1),(0,1,1),(1,0,1)\}.$
We have
$\operatorname{up}[Q]=\mathbb{N}^{3}\setminus\{(0,0,0)\}.$
Note that $(1,1,0)$ is essential in $\operatorname{up}[Q]$, so $(1,1,0)\in I_{\operatorname{up}[Q]}$ but $(1,1,0)\not\in Q$.
A poset $U\subseteq\mathbb{N}^{d}$ is an upset if $U=\operatorname{up}[S]$ for some set $S\subseteq\mathbb{N}^{d}$. Given an upset $U\subseteq\mathbb{N}^{d}$, define a monomial ideal
$\hat{U}=\langle x^{z}\mid z\in U\rangle\subseteq R^{d}.$
We omit the straightforward proof of the following:
Proposition 5.10.
The map $U\mapsto\hat{U}$ is a bijection between the set of upsets in $\mathbb{N}^{d}$ and the set of monomial ideals of $R^{d}$.
Lemma 5.11.
For $U\subseteq\mathbb{N}^{d}$ an upset and $z\in U$, let $L$ denote the full subposet of $\operatorname{dn}(z,U)$ with
$\operatorname{Ob}L=\{y\in\operatorname{dn}(z,U)\mid\|y-z\|_{1}\leq 2\}.$
The inclusion $L\hookrightarrow\operatorname{dn}(z,U)$ induces a bijection on path components.
Proof.
In what follows, we write $D=\operatorname{dn}(z,U)$. Since $U$ is an interval, each component of $D$ contains an element of the form $z-\mathbf{e}_{i}$ for some $\mathbf{e}_{i}$. Since $z-\mathbf{e}_{i}\in L$, the inclusion $L\hookrightarrow D$ induces a surjection on path components. It remains to show that this inclusion is an injection.
Consider $w,y\in L$ lying in the same component of $D$. We need to show that $w,y$ lie in the same component of $L$. There exists a path $w=q_{1},\ldots,q_{n}=y$ from $w$ to $y$ in $D$. We will modify this path to obtain a path from $w$ to $y$ lying in $L$. Clearly, we may assume $n>2$. We may assume without loss of generality that for each triple $q_{i-1},q_{i},q_{i+1}$ of consecutive points in the path, either
$q_{i-1}<q_{i}>q_{i+1}\quad\textup{or}\quad q_{i-1}>q_{i}<q_{i+1}.$
In the former case, we call $q_{i}$ an upper point, and in the latter case, we call $q_{i}$ a lower point.
If $w=z-\mathbf{e}_{j}-\mathbf{e}_{j^{\prime}}$ (for some $j$ and $j^{\prime}$), then by inserting a new element $w-\mathbf{e}_{j}$ in the path, if necessary, we may assume without loss of generality that $w<q_{2}$. Symmetrically, if $y=z-\mathbf{e}_{j}-\mathbf{e}_{j^{\prime}}$, then we may assume that $q_{n-1}>y$.
Since $U$ is an interval, we may further assume that each upper point $q_{i}$ is of the form $z-\mathbf{e}_{j}$, by replacing $q_{i}$ with a point above it, if necessary. Given these assumptions, for each lower point $q_{i}$, we have $q_{i-1}=z-\mathbf{e}_{j}$ and $q_{i+1}=z-\mathbf{e}_{j^{\prime}}$, so we have
$\wedge(q_{i-1},q_{i+1})=z-\mathbf{e}_{j}-\mathbf{e}_{j^{\prime}}\geq q_{i},$
where $\wedge$ denotes the meet. It follows that $\wedge(q_{i-1},q_{i+1})\in L$. Therefore, by replacing each lower point $q_{i}$ with $\wedge(q_{i-1},q_{i+1})$, we obtain a path from $w$ to $y$ in $L$. This shows that $w,y$ lie in the same component of $L$. Hence the map on components induced by the inclusion $L\hookrightarrow D$ is an injection, as desired. ∎
Lemma 5.12.
For $U\subseteq\mathbb{N}^{d}$ an upset, we have $E_{U}=\operatorname{Supp}\beta^{\hat{U}}_{1}$.
Proof.
By Corollary˜5.4, it suffices to check that $z\in E_{U}$ if and only if $|\pi_{0}(K^{z}(U))|\geq 2$. Note that $z\in E_{U}$ if and only if $\operatorname{dn}(z,U)$ has at least two components. Therefore, it suffices to exhibit a bijection between the components of $\operatorname{dn}(z,U)$ and those of $K^{z}(U)$. Lemma˜5.11 provides a bijection between the components of $\operatorname{dn}(z,U)$ and those of the subposet $L$ in the statement of that lemma. Moreover, components of $L$ correspond bijectively with components of the undirected graph $G$ underlying the Hasse diagram of $L$. Let $K^{+}$ denote the barycentric subdivision of the 1-skeleton of $K^{z}(U)$. The simplicial map $f\colon G\to K^{+}$ given by $f(z-\mathbf{e}_{s})=s$ is a graph isomorphism, and therefore induces a bijection on components. Composing these bijections yields a bijection between the components of $\operatorname{dn}(z,U)$ and those of $K^{z}(U)$. ∎
Lemma 5.13.
For $U\subseteq\mathbb{N}^{d}$ an upset, we have $I_{U}=\operatorname{Supp}\beta^{\hat{U}}_{0}\cup\operatorname{Supp}\beta^{\hat{U}}_{1}$.
Proof.
It is clear that $M_{U}=\operatorname{Supp}\beta^{\hat{U}}_{0}$ and by Lemma˜5.12, we have $E_{U}=\operatorname{Supp}\beta^{\hat{U}}_{1}$. Since $I_{U}=M_{U}\cup E_{U}$, the result follows. ∎
Example 5.14.
We identify a family of upsets $(U^{k})_{k\in\mathbb{N}}$ in $\mathbb{N}^{4}$ such that $U^{k}$ has $\Theta(k)$ minima and $|I_{U^{k}}|=\Theta(k^{2})$. By Lemma˜5.13, this corresponds to a family $(J^{k})_{k\in\mathbb{N}}$ of ideals of $R^{4}$ for which
$\sum_{z\in\mathbb{N}^{4}}\beta^{J^{k}}_{1}(z)=\Theta(|M_{J^{k}}|^{2}),$
as we needed in the proof of Corollary˜5.7.
Fixing $k\in\mathbb{N}$, let
$A$ $=\{(i,k-i,0,0)\mid i\in\{0,\ldots,k\}\},$
$B$ $=\{(0,0,i,k-i)\mid i\in\{0,\ldots,k\}\},$
$U^{k}$ $=\operatorname{up}[A\cup B].$
To simplify notation, we let $U=U^{k}$.
Note that $M_{U}=A\sqcup B$, so $|M_{U}|=2(k+1)$. We show that for each $i,j\in\{0,\dots,k\}$,
$t_{ij}\coloneqq(i,k-i,j,k-j)$
is essential. Indeed, writing
$D_{1}=\{(i,k-i,a,b)\mid(a,b)<(j,k-j)\},$
$D_{2}=\{(a,b,j,k-j)\mid(a,b)<(i,k-i)\},$
we have $\operatorname{dn}(t_{ij},U)=D_{1}\sqcup D_{2}$. Moreover, every point in $D_{1}$ is incomparable to every point in $D_{2}$. Any path in $\operatorname{dn}(t_{ij},U)$ connecting a point in $D_{1}$ to a point in $D_{2}$ must contain a pair of comparable points $p,p^{\prime}$, with $p\in D_{1}$ and $p^{\prime}\in D_{2}$. Thus, no such path can exist, which implies that $t_{ij}$ is essential. Therefore, $U$ has at least $(k+1)^{2}$ essential elements, so by Theorems˜5.5 and 5.13, we have $|I_{U}|=\Theta(k^{2})$, as claimed.
We are now ready to prove Theorem˜3.10. Recall the statement:
Theorem 3.10.
Let $Q$ be an interval in $\mathbb{N}^{d}$ with $n$ minima and let $P\subseteq Q$ be an initial scaffold of $Q$. We have
  • (i)
    $|P|=\Theta(n)$ for $d\leq 3$,
  • (ii)
    $|P|=\Theta(n^{2})$ for $d>3$.
Proof.
Suppose $P\subseteq Q$ is an initial scaffold of an interval $Q\subseteq\mathbb{N}^{d}$. Note that $\operatorname{dn}(p,Q)$ as at most $d$ components for each $p\in Q$. Thus, for each $p\in P$, there are at most $d$ relations in $P$ of the form $m<p$. Hence, to prove the result, it suffices to show that $|\operatorname{Ob}P|=|I_{Q}|$ satisfies the bounds of the theorem. By Lemma˜5.8, we have $|I_{Q}|\leq|I_{\operatorname{up}[Q]}|$. Since $\operatorname{up}[Q]\subseteq\mathbb{N}^{d}$ is an upset with the same minima as $Q$, Corollary˜5.7 and Lemma˜5.13 together imply that
$|I_{Q}|=\begin{cases}\Theta(n)&\textup{ for }d\leq 3,\\ \Theta(n^{2})&\textup{ for }d>3.\end{cases}$
The result follows. ∎
6. Computing Initial Scaffolds
6.1. Algorithm for Finite Posets
We next give an algorithm to compute the initial scaffold of a finite poset, thereby proving Theorem˜3.13. Recall the statement:
Theorem 3.13.
Given the Hasse diagram $(V,E)$ of a finite poset $Q$, we can compute an initial scaffold of $Q$ in time $O(|V||E|)$.
Proof.
If we choose an initial scaffold for each component of $Q$, then the union of these initial scaffolds is an initial scaffold for $Q$. The components of $Q$ can be computed in time $O(|V|+|E|)=O(|V||E|)$ by depth-first search on the Hasse diagram. It therefore suffices to treat the case of a connected poset $Q$. Assuming $Q$ is connected, the algorithm to compute an initial scaffold $P\subseteq Q$ then proceeds as follows: We refine the partial order on $Q$ to a total order via the standard topological sorting algorithm in time $O(|V|+|E|)=O(|E|)$. We process the elements of $Q$ in increasing order. When we visit $q\in Q$, we compute the components of $\operatorname{dn}(q,Q)$ using depth-first search in time $O(|E|)$. If the number of components is not 1, then we add $q$ to $P$ and, for each component $A$ of $\operatorname{dn}(q,Q)$, we add a relation $m<q$ in $P$ for some choice of minimum $m\in A$. These minima can be identified while computing the components of $\operatorname{dn}(q,Q)$ by counting in-degrees of vertices. It is clear that the total time cost of the algorithm is $O(|V||E|)$. ∎
6.2. Algorithm for Intervals in $\mathbb{N}^{3}$
The rest of this section is devoted to the proof of Theorem˜3.17. Let us recall the statement:
Theorem 3.17.
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=s$, we can compute an initial scaffold of $Q$ in time
  • (i)
    $O(s\log s)$ for $d=2,3$,
  • (ii)
    $O(s^{4})$ for $d>3$.
In the case that $Q$ is finite, these bounds also hold when $Q$ is instead specified by its extrema, with $s$ the total number of these.
We first prove Theorem˜3.17 (i) by giving an algorithm that computes an initial scaffold of an interval $Q\subseteq\mathbb{N}^{3}$ and bounding its complexity. The algorithm and its analysis specialize to intervals $Q\subseteq\mathbb{N}^{2}$ by embedding $\mathbb{N}^{2}$ into $\mathbb{N}^{3}$ via the map $(x,y)\mapsto(x,y,0)$.
Letting $U=\operatorname{up}[Q,\mathbb{N}^{3}]$, our algorithm first computes an initial scaffold $P$ of $U$ and then computes $P\cap Q$. According to the following result, this is an initial scaffold of $Q$.
Proposition 6.1.
If $Q\subseteq\mathbb{N}^{d}$ is an interval and $P$ is an initial scaffold of $\operatorname{up}[Q]$, then $P\cap Q$ is an initial scaffold of $Q$.
Proof.
This is immediate from Lemma˜5.8 and the definition of initial scaffold. ∎
The computation of $P$ proceeds by iterating through 2-D coordinate slices of $U$. To elaborate, for any upset $I\subseteq\mathbb{N}^{3}$ and $z\in\mathbb{N}$, the coordinate slice
$I_{z}\coloneqq\{(y_{1},y_{2})\in\mathbb{N}^{2}\mid(y_{1},y_{2},z)\in I\}$
is also an upset. To compute $P$, we iterate through values of $z\in\mathbb{N}$ in increasing order, maintaining the sets of minima $M_{U_{z}}$ sorted by $x$-coordinate. When we pass from level z1z-1 to level $z$, we update $M_{U_{z-1}}$ to obtain $M_{U_{z}}$ and record all elements of $E_{U}$ of the form $e=(x,y,z)$. For each such $e$, we compute the relations $m<e$ to be added to $P$.
The computation of $P\cap Q$ from $P$ is done in a similar but much simpler way, also by iterating through 2-D coordinate slices. As we explain below, whether the poset $Q$ is specified by its upset presentation or by its extrema, we use essentially the same algorithm to compute $P\cap Q$ from $P$.
We next explain in detail how we compute $P$. We begin by introducing some notation.
Notation 6.2.
  • For $m=(m_{1},m_{2})\in\mathbb{N}^{2}$ and $y\in\mathbb{N}$, we identify $(m,y)$ with $(m_{1},m_{2},y)\in\mathbb{N}^{3}$.
  • For $m\in M_{U_{z}}$, let $m^{\downarrow}=(m,y)\in\mathbb{N}^{3}$, where $y\in\mathbb{N}$ is the unique value such that $(m,y)\in M_{U}$.
  • Write $M_{U_{z}}$ in increasing order of $x$-coordinate as
    $M_{U_{z}}=(m_{1},\ldots,m_{k}).$
  • Let $W^{z}$ denote the set of elements of the form $(m_{i}\vee m_{i+1},z)$ such that for some $j\in\{i,i+1\}$, both of the following conditions hold:
    1. (1)
      $(m_{j},z)\in M_{U}$,
    2. (2)
      there is no $x\in M_{U_{z-1}}$ with $m_{j}<x<m_{i}\vee m_{i+1}$.
  • Let
    $X^{z}=\{(m,z)\mid m\in M_{U_{z-1}}\setminus M_{U_{z}}\}.$
Our approach to computing an initial scaffold $P$ of $U$ rests on the following proposition:
Proposition 6.3.
  • (i)
    $E_{U}=\bigcup_{z\in\mathbb{N}}(W^{z}\cup X^{z})$.
  • (ii)
    The relations of an initial scaffold $P\subseteq U$ are given as follows:
    1. (1)
      If $p=(m_{i}\vee m_{i+1},z)\in W^{z}$, then $P$ contains the relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$. (Note that either $m_{i}^{\downarrow}=(m_{i},z)$ or $m_{i+1}^{\downarrow}=(m_{i+1},z)$).
    2. (2)
      If $p=(m,z)\in X^{z}$, then $P$ contains the relation $m^{\downarrow}<p$. If in addition $p\not\in W^{z}$, then $P$ also contains exactly one relation $(a,z)<p$, where $(a,z)$ is an arbitrary element of $M_{U}$ with $a<m$.
We refer to the relations from (1) and (2) in Proposition˜6.3 (ii) as $W^{z}$-relations and $X^{z}$-relations, respectively.
Before proving Proposition˜6.3, we illustrate the result with an example.
Example 6.4.
Let
$U=\{(0,6,0),(1,5,0),(3,4,0),(4,2,0),(5,0,0),(1,3,1),(2,2,1),(4,1,1)\}.$
Then as illustrated in Fig.˜3, we have
$M_{U_{0}}$ $=\left((0,6),(1,5),(3,4),(4,2),(5,0)\right),$
$M_{U_{1}}$ $=\left((0,6),(1,3),(2,2),(4,1),(5,0)\right),$
$W^{0}$ $=\{((1,6),(3,5),(4,4),(5,2)\},$
$W^{1}$ $=\{(2,3,1),(4,2,1),(5,1,1)\},$
$X^{1}$ $=\{(1,5,1),(3,4,1),(4,2,1)\}.$
Note that $(4,2,1)\in W^{1}\cap X^{1}$. Also note that $(1,6,1)\not\in W^{1}$, because the second condition in the definition of $W^{z}$ is not satisfied.
xxyy12345123456
Figure 3. For the upset $U\subseteq\mathbb{N}^{3}$ of Example˜6.4, an illustration of $M_{U_{0}}$ (blue) and $M_{U_{1}}\setminus M_{U_{0}}$ (red), as well as the $(x,y)$-projections of $W^{0}$ (squares), $W^{1}$ (circles), and $X^{1}$ (diamonds).
Proposition˜6.3 (ii) yields an initial scaffold $P\subseteq U$ with
$\operatorname{Ob}P=M_{U_{0}}\cup M_{U_{1}}\cup W^{0}\cup W^{1}\cup X^{1}$
and the following relations:
$W^{0}$-relations:
$(0,6,0)$ $\leq(1,6,0)\geq(1,5,0)\leq(3,5,0)\geq(3,4,0)$
$\leq(4,4,0)\geq(4,2,0)\leq(5,2,0)\geq(5,0,0).$
$W^{1}$-relations:
$(1,3,1)\leq(2,3,1)\geq(2,2,1)\leq(4,2,1)\geq(4,1,1)\leq(5,1,1)\geq(5,0,0).$
$X^{1}$-relations:
$(1,5,0)$ $\leq(1,5,1)\geq(1,3,1),$
[…]
$(4,2,0)$ $\leq(4,2,1).$
Proof of Proposition˜6.3.
In what follows, all downsets are taken in $U$. Consider an element $p=(m_{i}\vee m_{i+1},z)\in W^{z}$, and let $j\in\{i,i+1\}$ be the index satisfying the condition in the definition of $W^{z}$ for $p$. Assume without loss of generality that $j=i$, so that $(m_{i},z)\in M_{U}$. To see that $p\in E_{U}$, it suffices to show that $(m_{i},z)$ and $(m_{i+1},z)$ belong to different components of $\operatorname{dn}(p)$. Clearly, $(m_{i},z),(m_{i+1},z)\in\operatorname{dn}(p)$. The component of $\operatorname{dn}(p)$ containing $m_{i}$ is
$\{(k,(m_{i})_{2},z)\mid(m_{i})_{1}\leq k<(m_{i+1})_{1}\}.$
This component does not contain $(m_{i+1},z)$. Hence $p$ is essential.
If $p=(m,z)\in X^{z}$, then $m\in M_{U_{z-1}}$. Moreover, there is some $a\in M_{U_{z}}$ with $a<m$. Note that $(m,z-1),(a,z)\in\operatorname{dn}(p)$. Since $m\in M_{U_{z-1}}$, it must be that $U$ contains no points of the form (b,z1)(b,z-1) with $b<m$. Thus, (m,z1)(m,z-1) and $(a,z)$ lie in different components of $\operatorname{dn}(p)$. Hence, $p$ is essential.
Conversely, suppose that $p=(p_{1},p_{2},z)$ is essential. Then since $U$ is an interval, at least two of the three vectors $\{p-\mathbf{e}_{1},p-\mathbf{e}_{2},p-\mathbf{e}_{3}\}$ lie in distinct components of $\operatorname{dn}(p)$. We first consider the case where $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ lie in distinct components of $\operatorname{dn}(p)$, which we denote as $C_{1}$ and $C_{2}$, respectively. We will show that in this case, $p\in W^{z}$. Since $C_{1}$ and $C_{2}$ are disjoint, we have $p-\mathbf{e}_{1}-\mathbf{e}_{2}\not\in\operatorname{dn}(p)$. Therefore,
$C_{1}$ $\subseteq\{(a,p_{2},c)\mid a\leq p_{1},\ c\leq z\},$
$C_{2}$ $\subseteq\{(p_{1},b,c)\mid b\leq p_{2},\ c\leq z\}.$
Note that $p=(m_{i}\vee m_{i+1},z)$ for some $i$; concretely,
$m_{i}$ $=\min\{(a,p_{2})\mid(a,p_{2},z)\in C_{1}\},$
$m_{i+1}$ $=\min\{(p_{1},b)\mid(p_{1},b,z)\in C_{2}\}.$
Moreover, since $C_{1}$ and $C_{2}$ are disjoint, they cannot both contain $p-\mathbf{e}_{3}$. Assume without loss of generality that $p-\mathbf{e}_{3}\not\in C_{1}$. We will show that $p$ satisfies the definition of an element of $W^{z}$, for $j=i$. Since $U$ is an interval, we have
$C_{1}\subseteq\{(a,p_{2},z)\mid a<p_{1}\}.$
This implies that $(m_{i},z)\in M_{U}$, and moreover that there exists no $x\in M_{U_{z-1}}$ with $m_{i}<x<m_{i}\vee m_{i+1}$. Hence $p\in W^{z}$, as desired.
Next, suppose that $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ do not lie in distinct components of $\operatorname{dn}(p)$, i.e., either $\{p-\mathbf{e}_{1},p-\mathbf{e}_{2}\}\not\subseteq\operatorname{dn}(p)$ or else $p-\mathbf{e}_{1}$ and $p-\mathbf{e}_{2}$ lie in the same component of $\operatorname{dn}(p)$ (in which case $p-\mathbf{e}_{1}-\mathbf{e}_{2}\in\operatorname{dn}(p)$). Assume without loss of generality that $p-\mathbf{e}_{1}\in\operatorname{dn}(p)$. Then $p-\mathbf{e}_{3}$ lies in a separate component of $\operatorname{dn}(p)$; call the respective components $U_{1}$ and $U_{3}$. Note that neither $p-\mathbf{e}_{3}-\mathbf{e}_{1}\in\operatorname{dn}(p)$ nor $p-\mathbf{e}_{3}-\mathbf{e}_{2}\in\operatorname{dn}(p)$, since otherwise we would have $U_{1}=U_{3}$, a contradiction. It follows that $p-\mathbf{e}_{3}\in M_{U_{z-1}}$, and since $p-\mathbf{e}_{1}\in U$, we also have that $p\not\in M_{U_{z}}$. Hence $p\in X^{z}$. This completes the proof of (i).
To prove (ii), suppose $p=(m_{i}\vee m_{i+1},z)\in W^{z}$. The proof of (i) shows that $(m_{i},z)$ and $(m_{i+1},z)$ lie in separate components of $\operatorname{dn}(p)$, so $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$ are valid choices of relations in $P$. If in addition, $p\not\in X^{z}$, then either $p-\mathbf{e}_{3}\not\in\operatorname{dn}(p)$ or $p-\mathbf{e}_{3}-e_{i}\in\operatorname{dn}(p)$ for some $j\in\{0,1\}$. In either case, $\operatorname{dn}(p)$ has two components, so no other relations with upper element $p$ need to be added to $P$.
Suppose $p=(m,z)\in X^{z}$, then since $p-\mathbf{e}_{3}\in\operatorname{dn}(p)$ and there is no $b<m$ with $(b,z-1)\in\operatorname{dn}(p)$, the interval property of $U$ implies that there is a component of $\operatorname{dn}(p)$ lying in the set $\{(m,y)\in\mathbb{N}^{3}\mid y<z$. It follows that $P$ must contain the relation $m^{\downarrow}<p$. If in addition $p\not\in W^{z}$, then $\operatorname{dn}(p)$ has exactly one component not containing $p-\mathbf{e}_{3}$, and all elements of this component have third coordinate $z$. Therefore, for $(a,z)$ an arbitrary element of $M_{U}$ with $a<m$, the relation $(a,z)<p$, is also valid in $P$, and since $\operatorname{dn}(p)$ has two components, no other relations with upper element $p$ need to be added to $P$.
If $p\in W^{z}\cap X^{z}$, then our treatment of the case $p\in W^{z}$ above yields two relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$, while the our treatment of the case $p\in X^{z}$ yields a third relation $m^{\downarrow}<p$. As $m_{i}^{\downarrow}$, $m_{i+1}^{\downarrow}$, and $m^{\downarrow}$ lie in different components of $\operatorname{dn}(p)$, all three relations can be included in $P$. Since $U$ is an interval in $\mathbb{N}^{3}$, $\operatorname{dn}(p)$ can have at most three components, so no other relations with upper element need to be added to $P$. ∎
Proof of Theorem˜3.17 (i).
We present an algorithm that iterates over $z$ to compute $M_{U_{z}}$, $W^{z}$ and $X^{z}$, as well as the relations $m<p$ of $P$ for $p\in W^{z}\cup X^{z}$.
Inductively, assume that we are given a list $A=M_{U_{z-1}}$, ordered by $x$-coordinate. This will ultimately be updated to $A=M_{U_{z}}$. Let
$B_{z}=\{m\in\mathbb{N}^{2}\mid(m,z)\in M_{U}\}.$
We initialize $X^{z}=W^{z}:=\{\}$ and an empty list $C:=\{\}$ that will contain deleted points from $A$ and is used in the computation of $W^{z}$. For each $m=(m_{x},m_{y})\in B_{z}$ we do the following: Let $L_{m}$ be the prefix of $A$ consisting of points with $y$-coordinate at least $m_{y}$, and let $R_{m}$ be the suffix of $A$ consisting of points with $x$-coordinate at least $m_{x}$. We then add all elements of $\{(n,z)\mid n\in L_{m}\cap R_{m}\}$ to $X^{z}$, along with the associated relations $n^{\downarrow}<(n,z)$. We also store the relation $(m,z)<(n,z)$ for possible inclusion into $P$; see below. We then update
$A\leftarrow(A\setminus(L_{m}\cap R_{m}))\cup\{m\},$
maintaining $A$ in sorted order by $x$-coordinate. The set of points $L_{m}\cap R_{m}$ that are deleted from $A$ are inserted into the list $C$, which is also maintained sorted in $x$-coordinate. Completing the iteration over all elements of $B_{z}$ produces $A=M_{U_{z}}$.
To compute $W^{z}$ and the remaining relations of $P$ whose upper element has third coordinate $z$, we proceed as follows: Writing
$M_{U_{z}}=(m_{1},\ldots,m_{k})$
as above, for each $p=m_{i}\vee m_{i+1}$ where at least one of $m_{i}$ or $m_{i+1}$ is in $M_{U}$, we check whether $p\in W^{z}$. Suppose, as in the definition of $W^{z}$, that $j\in\{i,i+1\}$. Then provided each $m_{j}$ is stored together with the third coordinate of $m_{j}^{\downarrow}$, checking whether $(m_{j},z)\in M_{U}$ is trivial. Checking whether $\nexists\ x\in M_{U_{z-1}}$ with $m_{j}<x<m_{i}\vee m_{i+1}$ amounts to locating $m_{j}$ in the sorted list of points that are deleted from $A$ while computing $M_{U_{z}}$. So, we locate $m_{j}$ in the sorted list $C$ through a binary search and then check if the point in $C$ immediately to the right of $m_{j}$ (the point in $C$ above $m_{j}$ with the least $x$-coordinate if it exists) has the same $y$-coordinate as $m_{j}$. Such a point exists if and only if the second condition in the definition of $W^{z}$ is satisfied for the given choice of $j$. If $m_{i}\vee m_{i+1}\in W^{z}$, then we include the relations $m_{i}^{\downarrow}<p$ and $m_{i+1}^{\downarrow}<p$. Once $W^{z}$ has been computed, for each $(n,z)\in X^{z}\setminus W^{z}$, we insert the relation $(m,z)<(n,z)$ computed earlier into $P$.
To explain how we compute $P\cap Q$ from $P$, assume first that $Q$ is specified as input to our algorithm by its upset presentation $(M_{Q},M_{Q^{\prime}})$. We sweep through the 2-D slices $Q^{\prime}_{z}$} of $Q^{\prime}$, computing each $M_{Q^{\prime}_{z}}$ from $M_{Q^{\prime}_{z-1}}$ exactly as we computed $U_{z}$ from $U_{z-1}$ above. For $e=(x,y,z)\in E_{U}$, note that $e\not\in Q$ if and only if $e\in Q^{\prime}$ if and only if $(x,y)\geq m$ for some $m\in M_{Q^{\prime}_{z}}$. The last condition can be checked using at most two binary searches on $M_{Q^{\prime}_{z}}$,
If $Q$ is instead specified by its extrema, a slight variant of the above procedure computes $P\cap Q$ from $P$: Let $U^{\prime}$ denote the upset of the set of maxima of $Q$. We sweep through the 2-D slices $U^{\prime}_{z}$} of $U^{\prime}$, computing each $M_{U^{\prime}_{z}}$ from $M_{U^{\prime}_{z-1}}$. Given $e=(x,y,z)\in E_{U}$ and $M_{U^{\prime}_{z-1}}$, note that $e\in Q$ if and only if $(x,y)\leq m$ for some $m\in M_{U^{\prime}_{z}}$. As above, the second condition can be checked by at most two binary searches on $M_{U^{\prime}_{z}}$,
This completes the specification of our algorithm. We now turn to the complexity analysis. For each $m\in B_{z}$, performing the update of the list $A$ requires time $O(\log|A|)=O(\log|M_{U}|)$; the time cost is dominated by a pair of binary searches in $A$ to find the (possibly null) minimum elements of $L_{m}$ and $R_{m}$. Updating $A$ requires deleting elements of $L_{m}\cap R_{m}$ which can be done in time $O(|L_{m}\cap R_{m}|)$. Thus, this update of $A$ takes $O(\sum_{m\in B}|L_{m}\cap R_{m}|)$. Since each element in $M_{U}$ gets inserted and then possibly deleted only once from $A$ over all levels of $z$, the total update time for $A$ is
$O(\sum_{z}\sum_{m\in B_{z}}|L_{m}\cap R_{m}|)=O(|M_{U}|).$
Similarly, binary searches across all levels of $z$ occur $O(1)$ times for every element in $M_{U}$. Thus, we have $O(|M_{U}|\log|M_{U}|)$ time for all binary searches. Therefore, computing $M_{U_{z}}$ for all values of $z$ takes time
$O(|M_{U}|+|M_{U}|\log|M_{U}|)=O(|M_{U}|\log|M_{U}|)=O(s\log s).$
Computation of $X^{z}$ is dominated by the time to determine the set $L_{m}\cap R_{m}$ for each $m\in B_{z}$. Hence, again this computation takes $O(|M_{U}|\log|M_{U}|)=O(s\log s)$ time over all levels of $z$. Notice that, we check only those points $m_{i}\vee m_{i+1}$ for inclusion in $W^{z}$ where at least one of $m_{i}$ or $m_{i+1}$ is in $M_{U}$. Thus, over all $z$, we check at most $O(|M_{U}|)$ points for inclusion into the set $W^{z}$. Each check for such an inclusion takes only $O(1)$ binary searches in the list $C$. This list can have at most $O(|M_{U}|)$ elements because $C$ contains only elements in $M_{U}$ each of which is deleted from $A$ only once. Thus, computation of $W^{z}$ over all levels of $z$ takes time
$O(|M_{U}|\log|M_{U}|)=O(s\log s).$
To compute $P\cap Q$, we must compute either $M_{Q^{\prime}_{z}}$ or $M_{U^{\prime}_{z}}$ at all values of $z$, which takes time $O(s\log s)$. Besides this, we must perform $O(|E_{U}|)$ binary searches on lists of size $O(s)$. By Theorem˜3.10 (i), we have
$|E_{U}|=O(|M_{U}|)=O(|M_{Q}|)=O(s),$
so the total time required by all of these binary searches is $O(s\log s)$. Thus, the total time to compute $P\cap Q$ from $P$ is $O(s\log s)$. The result follows. ∎
6.3. Algorithm for Intervals in $\mathbb{N}^{d}$
We next prove Theorem˜3.17 (ii) by giving an algorithm to compute an initial scaffold of an interval in $\mathbb{N}^{d}$ for arbitrary $d$, and bounding its complexity. As in the $d=3$ case, our strategy is to first compute an initial scaffold $P$ of $U=\operatorname{up}[Q,\mathbb{N}^{d}]$ and then compute the intersection $P\cap Q$, which is a hull skeleton of $Q$ by Proposition˜6.1.
To compute $P$, we use following structural result:
Proposition 6.5.
For $T\subseteq U$ the full subposet consisting of elements of $M_{U}$ and their pairwise joins, $P\subset U$ is an initial scaffold of $U$ if and only if it is an initial scaffold of $T$.
Proof.
It is easily checked that any element of $E_{U}$ is a pairwise join of elements in $M_{U}$. (Equivalently, any element in the support of $\beta_{1}^{\hat{U}}$ is such a join, which is immediate by considering the Taylor resolution of $\hat{U}$ [miller2005combinatorial].) Hence $E_{U}\subseteq T$. We claim that for any $t\in T$, the inclusion $\operatorname{dn}(t,T)\hookrightarrow\operatorname{dn}(t,U)$ induces a bijection on path components. The claim implies that $I_{T}=I_{U}$. Since $M_{T}=M_{U}$, the claim then also implies that $T$ and $U$ have the same initial scaffolds, as desired.
It remains to prove the claim. Since $M_{T}=M_{U}$, the inclusion $\operatorname{dn}(t,T)\hookrightarrow\operatorname{dn}(t,U)$ clearly induces a surjection on path components. We need to show that the induced map is also an injection. Suppose $w,y\in T$, with $w$ and $y$ contained in the same component of $\operatorname{dn}(t,U)$. It suffices to show that $w$ and $y$ are contained in the same component of $\operatorname{dn}(t,T)$. Our argument is reminiscent of the proof of Lemma˜5.11. By assumption, there exists a path $w=q_{1},\ldots,q_{n}=y$ from $w$ to $y$ in $\operatorname{dn}(t,U)$. We may assume without loss of generality that for each triple $q_{i-1},q_{i},q_{i+1}$ of consecutive points in the path, either $q_{i-1}\leq q_{i}\leq q_{i+1}$ or $q_{i-1}\geq q_{i}\leq q_{i+1}$. In the former case, we call $q_{i}$ an upper point, and in the latter case, we call $q_{i}$ a lower point. Moreover, we may assume without loss of generality that $n>2$, and that $q_{2}$ and $q_{n-1}$ are lower points. Replacing each lower point with a minimum below it, and then replacing each upper point $q_{i}$ with the join $q_{i-1}\vee q_{i+1}$, we obtain a path from $w$ to $y$ in $\operatorname{dn}(t,T)$. Thus, $w$ and $y$ lie in the same component of $\operatorname{dn}(t,T)$, as we wanted. ∎
Proof of Theorem˜3.17 (ii).
By Proposition˜6.5, to compute an initial scaffold of $U=\operatorname{up}[Q]$, it suffices to compute an initial scaffold of $T$. To do this, we first compute $T\setminus M_{T}$ and then order this set in a way compatible with the partial order induced by $\mathbb{N}^{d}$, e.g., lexicographically. We then iterate through the elements of $T\setminus M_{T}$ in order, adding each $t\in E_{T}$ to $P$, together with the required relations with upper element $t$. To elaborate, let $P^{t}$ be the portion of $P$ computed so far when we come to $t\in T\setminus M_{T}$. We compute $P^{t}\cap\operatorname{dn}(t,T)$ by comparing each element of $P^{t}$ to $t$. Similar to our algorithm for computing initial scaffolds of general finite posets given in Theorem˜3.13, we then compute the components of $P^{t}\cap\operatorname{dn}(t,T)$ via depth-first search. Proposition˜4.7 implies that it suffices here to consider the components of $P^{t}\cap\operatorname{dn}(t,T)$ rather than the components of $\operatorname{dn}(t,T)$. If $P^{t}\cap\operatorname{dn}(t,T)$ has more than one component, then we add $t$ to $P$, and for each component $A$ of $P^{t}\cap\operatorname{dn}(t,T)$, we chose a minimum $m\in A$ and add the relation $m<t$ to $P$.
For each element $t\in T\setminus M_{T}$ considered in the above algorithm, identifying the components of $P^{t}$ requires time $O(|P|)=O(|M_{Q}|^{2})=O(s^{2})$. At most $d$ edges are added to $P^{t}$ per element $t$, so adding these edges requires constant time. Since $T$ has size $O(|M_{Q}|^{2})=O(s^{2})$, the total time to compute an initial scaffold of $P$ is $O(s^{4})$.
Given $P$, we can straightforwardly compute $P\cap Q$ in time $O(s^{3})$. To explain, if the input to our algorithm is the upset presentation $(M_{Q},M_{Q^{\prime}})$, then for each $e\in E_{U}$, checking whether $e\in P$ amounts to checking whether $e\geq m$ for some $m\in M_{Q^{\prime}}$, as we have seen in the proof of Theorem˜3.17 (i). Since $|M_{Q^{\prime}}|=O(s)$ and $|E_{U}|=O(s^{2})$ by Theorem˜3.10 (i), the total time required for all such comparisons is $O(s\cdot s^{2})=O(s^{3})$. If $Q$ is instead given by the minima and the maxima of $Q$, the computation of $P\cap Q$ is analogous. Thus, the total time to compute an initial scaffold of $Q$ is $O(s^{4})$. ∎
7. Proofs of Results on Limit Computation
In this section, we complete the proofs of our results on limit computation by proving Proposition˜3.26 and Corollary˜3.31 (i).
7.1. Proof of Proposition˜3.26
Recall the statement of Proposition˜3.26:
Proposition 3.26.
Suppose we are given
  1. (1)
    a poset $Q$, represented in such a way that pairs of elements can be compared in constant time,
  2. (2)
    a $(Q,r)$-complex with homology $H$, and
  3. (3)
    the set of relations of a subposet $i\colon P\hookrightarrow Q$.
We can compute a matrix representation of $H\circ i$ in time $O(|P|r^{\omega})$.
We prepare for the proof by recalling some results from elementary linear algebra, which we state without proof.
Definition 7.1.
For $\mathbbm{k}$-vector spaces $U\subseteq V$, we say a set of vectors $C$ in $V$ is a basis extension for $U\subseteq V$ if for some (or, equivalently, any) basis $B$ for $U$, $B\sqcup C$ is a basis for $V$.
Lemma 7.2.
If $C$ is a basis extension for $U\subseteq V$, then the quotient map $V\to V/U$ sends $C$ to a basis $\bar{C}$ for $V/U$.
For a linear map $\alpha\colon V\to V^{\prime}$ and ordered bases $B$ and $B^{\prime}$ for $V$ and $V^{\prime}$, we write the matrix representation of $\alpha$ with respect to these bases either as $\alpha]^{B^{\prime},B}$, or more simply as $\alpha]$.
Lemma 7.3.
Given a linear map $\alpha\colon V\to V^{\prime}$ with $\alpha(U)\subseteq U^{\prime}$, let $\bar{\alpha}\colon V/U\to V^{\prime}/U^{\prime}$ denote the induced map. Let $B,B^{\prime}$ be ordered bases for $U,U^{\prime}$, and let $C,C^{\prime}$ be ordered basis extensions for $U\subseteq V$, $U^{\prime}\subseteq V^{\prime}$. Then the matrix $\bar{\alpha}]^{\bar{C}^{\prime},\bar{C}}$ is the block of the matrix $\alpha]^{B^{\prime}\sqcup C^{\prime},B\sqcup C}$ consisting of columns indexed by $C$ and rows indexed by $C^{\prime}$.
Definition 7.4.
  • (i)
    Given a non-zero vector $c=(c_{1},\ldots,c_{l})\in\mathbbm{k}^{l}$, the pivot index of $c$ is the smallest index $k$ such that $c_{k}\neq 0$.
  • (ii)
    Given a finite dimensional vector space $V$ with ordered basis $A$ and $v\in V$, let $v]\in\mathbbm{k}^{|A|}$ be the representation of $v$ with respect to $A$. We say that an ordered set of vectors $C=\{c_{1},\ldots,c_{l}\}$ in $V$ is in echelon form (with respect to $A$) if the pivot indices of the $c_{k}]$ are strictly increasing with $k$.
Proof of Proposition˜3.26.
Recalling Definition˜3.25, suppose that the given matrices $[f]$ and $[g]$ represent $f\colon X\to Y$ and $g\colon Y\to Z$ with respect to ordered bases $A^{X},A^{Y},A^{Z}$ for $X,Y,Z$. For each $p\in P$, these bases induce ordered bases $A^{X}_{p},A^{Y}_{p},A^{Z}_{p}$ for $X_{p},Y_{p},Z_{p}$. The matrix representation $f_{p}]$ of $f_{p}$ with respect to the induced bases is obtained from $[f]$ by taking the submatrix consisting of all rows and columns with labels $\leq p$, and similarly for $g_{p}]$. Since we assume that pairs of elements of $Q$ can be compared in constant time, deciding whether a column or row of $[f]$ belongs to $f_{p}]$ also takes constant time.
Our algorithm works as follows: First, for each $p\in P$, we use Gaussian elimination with backsolve to compute a basis $B_{p}$ for $\operatorname{im}f_{p}$, a basis extension $C_{p}$ for $\operatorname{im}f_{p}\subseteq\ker g_{p}$, and a basis extension $D_{p}$ for $\ker g_{p}\subseteq Y_{p}$, where all elements are given as linear combinations of the basis $A^{Y}_{p}$. Our algorithm selects the basis $\bar{C}_{p}$ for $H_{p}$.
We next explain the computation of $B_{p}$, $C_{p}$, and $D_{p}$ in detail. To begin, we compute an ordered basis $A^{\ker g}_{p}$ of $\ker g_{p}$ in echelon form (with respect to $A^{Y}_{p}$). We do this by applying Gaussian elimination with backsolve to $g_{p}]$. Writing $A^{Y}_{p}=\{a_{1},\ldots,a_{l}\}$, we take
[…]
Multiplying $f]=[f]^{A^{Y}_{p},A^{X}_{p}}$ by a change-of-basis matrix on the left, we obtain the matrix $f]^{(A^{\ker g}_{p}\sqcup D_{p}),A^{X}_{p}}$. Let $\gamma_{p}\colon X_{p}\to\ker g_{p}$ denote the co-restriction of $f_{p}$ to $\ker g_{p}\subseteq Y_{p}$, i.e., $\gamma_{p}(v)=f_{p}(v)$ for all $v\in X_{p}$. We obtain the matrix $\gamma_{p}]^{A^{\ker g}_{p},A^{X}_{p}}$ as the top block of $[f]$. Performing Gaussian elimination on the transpose of this matrix yields a basis $B_{p}$ for $\operatorname{im}\gamma_{p}$ in echelon form with respect to $A^{\ker g}_{p}$. We now compute the basis extension $C_{p}$ for $\operatorname{im}f_{p}\subseteq\ker g_{p}$ in the same way we computed $D_{p}$. By the way $C_{p}$ is constructed, we have $C_{p}\subseteq A^{\ker g}_{p}$, so we immediately have expressions for the elements of $C_{p}$ in the basis $A^{Y}_{p}$. To express the elements of the basis $B_{p}$ in the basis $A^{Y}_{p}$, it suffices to perform a single matrix multiplication.
For each relation $p\leq q$ in $P$, we must compute the matrix $H_{pq}]^{\bar{C}_{q},\bar{C}_{p}}$. To do this, first we form the matrix
$Y_{pq}]^{A^{Y}_{q},A^{Y}_{p}}=\begin{pmatrix}I\\ 0\end{pmatrix}$.
We then multiply this matrix by change-of-basis matrices on both the left and right to obtain the matrix
$W\coloneqq[Y_{pq}]^{(B_{q}\sqcup C_{q}\sqcup D_{q}),(B_{p}\sqcup C_{p}\sqcup D_{p})}.$
It follows from Lemma˜7.3 that $H_{pq}]^{\bar{C}_{q},\bar{C}_{p}}$ is the block of $W$ with columns indexed by $C_{p}$ and rows indexed by $C_{q}$.
It remains to show that this algorithm runs in time $O(|P|r^{\omega})$. For each element or relation of $P$, the algorithm performs Gaussian elimination with backsolve a constant number of times and also performs a constant number of matrix multiplications. Each matrix considered has dimension $O(r)\times O(r)$, so each instance of Gaussian elimination or matrix multiplication requires time $O(r^{\omega})$; see Proposition˜3.21. Since such operations dominate the cost of the algorithm, it follows that our algorithm runs in time $O(|P|r^{\omega})$. ∎
7.2. Proof of Corollary˜3.31 (i)
Let us recall the statement of Corollary˜3.31 (i), i.e., the case $d=2$ of the corollary:
Corollary 3.31 (i).
Given an $(\mathbb{N}^{2},r)$-complex with homology $H$ and the upset presentation of an interval $i\colon Q\hookrightarrow\mathbb{N}^{2}$ with $s=\|Q\|$, we can compute $\lim(H\circ i)$ in time $O(s\log s+r^{3})$.
Proof.
A standard result in quiver representation theory [gabriel1972unzerlegbare, carlsson2010zigzag] says that for $P$ a finite zigzag poset (see Example˜3.4 and a functor $G\colon P\to\mathbf{Vec}$, there is a unique multiset $\mathcal{B}(G)$ of intervals in $P$ such that
(7.1) $G\cong\bigoplus_{I\in\mathcal{B}(G)}\mathbbm{k}^{I}$
where the $\mathbbm{k}^{I}$ are the interval modules of Definition˜2.14. If all vector spaces of $G$ have finite dimension, then $\mathcal{B}(G)$ is finite, so the direct sum decomposition of $G$ above is in fact a direct product. Since small limits and categorical products commute [riehl2017category, Theorem 3.8.1], we have
$\lim G\cong\bigoplus_{I\in\mathcal{B}(G)}\lim\mathbbm{k}^{I}.$
Moreover, it is easily checked that for each interval $I\subseteq P$, we have
$\lim\mathbbm{k}^{I}=\begin{cases}\mathbbm{k}&\textup{ if $I$ is a downset},\\ 0&\textup{otherwise.}\end{cases}$
It follows that we can compute $\lim G$ by computing a decomposition of $G$ into indecomposable modules $\mathbbm{k}^{I}$ as in Eq.˜7.1. However, an important subtlety here is that, to compute a limit cone for $G$ (and not only for some functor naturally isomorphic to $G$), it is not enough to compute $\mathcal{B}(G)$; we also need an explicit isomorphism
(7.2) $\bigoplus_{I\in\mathcal{B}(G)}\mathbbm{k}^{I}\xrightarrow{\cong}G.$
We now apply these ideas to our setting. By Example˜3.4, $Q$ has a unique initial scaffold $j\colon P\hookrightarrow Q$, where $P$ is a finite zigzag poset. Let $G$ be the restriction of $H$ to $P$, i.e., $G=H\circ i\circ j$. We can compute $\mathcal{B}(G)$ and the isomorphism of Eq.˜7.2 in time $O(r^{3})$ by applying the zigzag persistence algorithm of [DTM25] to the restriction of the chain complex $\operatorname{\mathcal{C}}_{\mathbb{N}^{2}}^{r}$ to $P$. By Theorem˜3.17 (i), we can compute $P$ in time $O(s\log s)$. These two computations dominate the cost of computing $\lim(G)$. Hence, this approach computes $\lim G$, and thus $\lim(H\circ i)$, in total time $O(s\log s+r^{3})$. ∎
Remark 7.5.
If we wish to only compute $\dim(\lim(G\circ i))$ or the limit cone of a functor naturally isomorphic to $G\circ i$, then the complexity bound of Corollary˜3.31 (i) can be improved $O(s\log s+r^{\omega})$ by using either the zigzag persistence algorithm of [milosavljevic2011zigzag] or the more practical algorithm of [DH22].
8. Proofs of Results on Generalized Rank Computation
In this section, we prove Corollary˜3.36 and Corollary˜3.37, our bounds on the cost of generalized rank computation. We use the following notation:
Notation 8.1.
Given a functor $G\colon Q\to\mathbf{Vec}$ where $Q$ is a finite, connected poset, let $j^{\mathcal{I}}\colon P^{\mathcal{I}}\hookrightarrow Q$ and $j^{\mathcal{F}}\colon P^{\mathcal{F}}\hookrightarrow Q$ be initial and final scaffolds of $Q$. Given a choice of minimum $m\in Q$ and a maximum $w\in Q$ with $m\leq w$, let $\{m\leq w\}$ denote the two-element poset with a single non-identity relation $m\leq w$. Let
$P=P^{\mathcal{I}}\cup P^{\mathcal{F}}\cup\{m\leq w\}$
and let $j\colon P\hookrightarrow Q$ be the inclusion.
We first prove Corollary˜3.36. Recall the statement:
Corollary 3.36.
Let $Q$ be a finite, connected poset with $s$ extrema and let $^{\mathcal{I}}, PP^{\mathcal{F}}$ denote initial and final scaffolds of $Q$. Given the Hasse diagram $(V,E)$ of $Q$ and a $(Q,r)$-complex with homology $H$, we can compute $\operatorname{grank}(H)$ in time
$O(|V||E|+(|P^{\mathcal{I}}|+|P^{\mathcal{F}}|)s^{\omega-1}r^{\omega}).$
Proof.
To compute $\operatorname{grank}(H)$, we first compute initial and final scaffolds $j^{\mathcal{I}}\colon P^{\mathcal{I}}\hookrightarrow Q$ and $j^{\mathcal{F}}\colon P^{\mathcal{F}}\hookrightarrow Q$ via Theorem˜3.13. Choosing $m\leq w$ as in ˜8.1 and letting $j\colon P\hookrightarrow Q$ be as defined there, we compute a matrix representation of $H\circ j$ via Propositions˜3.26 and 3.29, which, in particular, yields a basis $B_{p}$ of $H_{p}$ for each $p\in P$. This restricts to matrix representations of $H\circ j^{\mathcal{I}}$ and $H\circ j^{\mathcal{F}}$, which we use to compute $\lim(H\circ j^{\mathcal{I}})$ and $\operatorname{colim}(H\circ j^{\mathcal{F}})$ as in Corollary˜3.30.
Let $\gamma_{m}\colon\lim H\to H_{m}$ and $\gamma^{\prime}_{w}\colon H_{w}\to\operatorname{colim}H$ be the respective cone and cocone maps. Observe that
$\operatorname{grank}H=\operatorname{rank}(\gamma^{\prime}_{w}\circ H_{mw}\circ\gamma_{m}).$
Since $m$ is a minimum, a matrix representation $\gamma_{m}]$ of $\gamma_{m}$ with respect to the basis $B_{m}$ is given by coordinate projection of the computed presection basis of $\lim H$; see ˜3.19 (i). Dually, coordinate projection also yields a matrix representation $\gamma^{\prime}_{w}]$ of $\gamma^{\prime}_{w}$ with respect to the basis $B_{w}$. The matrix representation of $H\circ j$ furnishes a matrix representation of $H_{mw}]$ with respect to the bases $B_{m}$ and $B_{w}$. To compute $\operatorname{grank}H$, we compute the matrix product $\gamma^{\prime}_{w}][H_{mw}][\gamma_{m}]$ and perform Gaussian elimination on the result.
The cost of these computations is dominated by the cost of computing $\lim H$ and $\operatorname{colim}H$, which is given by Corollary˜3.30. The result follows. ∎
We next turn to the proof of Corollary˜3.37. Recall the statement:
Corollary 3.37.
Given an $(\mathbb{N}^{d},r)$-complex with homology $H$ and the $s$ extrema of a finite interval $i\colon Q\hookrightarrow\mathbb{N}^{d}$, we can compute $\operatorname{grank}(H\circ i)$ in time
  • (i)
    $O(s\log s+r^{\omega})$ for $d=2$,
  • (ii)
    $O((rs)^{\omega})$ for $d=3$,
  • (iii)
    $O(r^{\omega}s^{\omega+1})$ for $d>3$.
Proof of Corollary˜3.37 (ii,iii).
The proof is the same as the proof of Corollary˜3.36, using Theorems˜3.17, 3.31 and 3.32 in place of Theorems˜3.13 and 3.30. ∎
In the setting of Corollary˜3.37 (i), the proof of Corollary˜3.36, together with Corollary˜3.31 (i), yields the bound $O(s\log s+r^{3})$ for generalized rank computation over a finite interval in $\mathbb{N}^{2}$, but does not yield the stronger bound $O(s\log s+r^{\omega})$ given by Corollary˜3.37 (i). The term $r^{3}$ arises from the cost of (co)limit computation via Corollary˜3.31 (i); see Remark˜7.5. We therefore need a different argument to prove Corollary˜3.37 (i). Our proof hinges on following lemma.
Proposition 8.2.
In the setting of ˜8.1, we have
$\operatorname{grank}(G)=\operatorname{grank}(G\circ j).$
We will only need the case of Proposition˜8.2 where $Q$ is an interval in $\mathbb{N}^{2}$, but it is no more difficult to prove the general result. As noted in Remark˜3.7, a variant of the proposition was obtained independently in [brustle2025generalized].
Proof of Proposition˜8.2.
Let $k^{\mathcal{I}}:P^{\mathcal{I}}\hookrightarrow P$ and $k^{\mathcal{F}}:P^{\mathcal{F}}\hookrightarrow P$ be the inclusions. Note that $j^{\mathcal{I}}=j\circ k^{\mathcal{I}}$ and $j^{\mathcal{F}}=j\circ k^{\mathcal{F}}$. By Theorem˜3.5, $j^{\mathcal{I}}$ is initial, so it induces an isomorphism
$\lim G\to\lim(G\circ j^{\mathcal{I}})=\lim(G\circ j\circ k^{\mathcal{I}}).$
By the functoriality of limits with respect to the index category, this isomorphism factors as
(8.1) $\lim G\hookrightarrow\lim(G\circ j)\twoheadrightarrow\lim(G\circ j^{\mathcal{I}}),$
where the first map is an injection, the second map is a surjection, and both maps commute with the relevant cone maps.
Dually, the map $j^{\mathcal{F}}$ induces an isomorphism $\operatorname{colim}(G\circ j^{\mathcal{F}})\to\operatorname{colim}G$ and by functoriality, this map factors as
(8.2) $\operatorname{colim}(G\circ j^{\mathcal{F}})\hookrightarrow\operatorname{colim}(G\circ j)\twoheadrightarrow\operatorname{colim}G,$
where again, the first map is an injection and the second map is an surjection. Let $\alpha$ denote the composite map
$\lim(G\circ j)\to G_{m}\to G_{w}\to\operatorname{colim}(G\circ j)$
and let $\beta$ denote the composite map
$\lim(G\circ j^{\mathcal{I}})\to G_{m}\to G_{w}\to\operatorname{colim}(G\circ j^{\mathcal{F}}).$
Note that $\operatorname{grank}(G\circ j)=\operatorname{rank}(\alpha)$. Moreover, since $j^{\mathcal{I}}$ and $j^{\mathcal{F}}$ are initial and final, respectively, we also have $\operatorname{grank}(G)=\operatorname{rank}\beta$. The map $\alpha$ factors as
$\lim(G\circ j)\twoheadrightarrow\underbrace{\lim(G\circ j^{\mathcal{I}})\to G_{m}\to G_{w}\to\operatorname{colim}(G\circ j^{\mathcal{F}})}_{\beta}\hookrightarrow\operatorname{colim}(G\circ j),$
where the first factor is the surjection from Eq.˜8.1 and the last factor is the injection from Eq.˜8.2.
Observe that for any linear maps $f\colon V^{\prime}\to V$, $g\colon V\to W$, and $h\colon W\to W^{\prime}$ with $f$ a surjection and $h$ an injection, we have $\operatorname{rank}(h\circ g\circ f)=\operatorname{rank}g$. Applying this to the maps above yields
$\operatorname{grank}(G\circ j)=\operatorname{rank}(\alpha)=\operatorname{rank}(\beta)=\operatorname{grank}(G).\qed$
Proof of Corollary˜3.37 (i).
We use an argument similar to the proof of Theorem˜3.35 in [DKM22], but framed in our general formalism of initial and final scaffolds. We compute initial and final scaffolds $P^{\mathcal{I}}$ and $P^{\mathcal{F}}$ of $Q$ in time $O(s\log s)$ using Theorem˜3.17 (i). Recall from Example˜3.4 that $P^{\mathcal{I}}$ is a zigzag poset; by duality, the same is true for $P^{\mathcal{F}}$. Let $m$ and $w$ be the leftmost minimum and maximum of $Q$, respectively. Then $m\in P^{\mathcal{I}}$, $w\in P^{\mathcal{F}}$, and $m\leq w$. Let $P$ and $j$ be as in ˜8.1 and let $G=H\circ i$. Then $P$ is a zigzag poset and the barcode of the zigzag persistence module $G\circ j$ can be computed in time $O(s\log s+r^{\omega})$ via the algorithm of either [milosavljevic2011zigzag] or [DH22]. The number of copies of the interval $P$ in the barcode is the number of interval summands of $G\circ j$ with support $P$, which in turn is $\operatorname{grank}(G\circ j)$ by [chambers2018persistent, Corollary 3.2]. We have $\operatorname{grank}(G\circ j)=\operatorname{grank}(G)=\operatorname{grank}(H\circ i)$ by Proposition˜8.2. The result follows. ∎
9. Questions for Future Work
We imagine that it may be possible to fully describe and efficiently compute minimal initial functors valued in arbitrary very small categories, extending our results about posets. It would also be worthwhile to study the optimality of our algorithmic results, particularly for limit computation. Finally, given that limits and diagrams of vectors spaces indexed by posets are both ubiquitous in mathematics, we are hopeful that the results of this paper could find practical use in applications beyond the generalized rank computations considered here. The development of such applications remains a subject for future work.